### v1

# PQA: Zero-shot Protein Question Answering for Free-form Scientific Enquiry with Large Language Models

Abstract
We introduce the novel task of zero-shot Protein Question Answering (PQA) for free-form scientific enquiry. Given a previously unseen protein sequence and a natural language question, the task is to deliver a scientifically accurate answer. This task not only supports future biological research, but could also provide a test bed for assessing the scientific precision of large language models (LLMs). We contribute the first specialized dataset for PQA model training, containing 257K protein sequences annotated with 1.97M scientific question-answer pairs. Additionally, we propose and study several novel biologically relevant benchmarks for scientific PQA. Employing two robust multi-modal architectures, we establish an initial state-of-the-art performance for PQA and reveal key performance factors through ablation studies. Our comprehensive PQA framework, named Pika, including dataset, code, model checkpoints, and a user-friendly demo, is openly accessible on github.com/EMCarrami/Pika, promoting wider research and application in the field.
Machine Learning, Multimodel Large Language Models, Protein Sequence
11footnotetext: Relation Therapeutics, London, UK22footnotetext: Ludwig Maximilian University of Munich, Germany*Corresponding author: Eli M Carrami, eli@relationrx.com, eli.carrami@gmail.com
1 Introduction
Figure 1: Schematic of Pika framework. Pika-DS is created from filtered SwissProt entries followed by processing using GPT3.5.
Proteins, essential to biological functions, are complex macromolecules that perform a myriad of cellular roles determined by their complex structures and interactions. As a polymer of amino acids, drawn from a pool of 20 natural ones, the sequence in which these building blocks are arranged dictates the protein’s three-dimensional structure, which is critical for its function. These structures facilitate interactions with various molecules, underpinning numerous biological processes. Given the critical role of proteins in both fundamental biology and applied biomedical research, a deep understanding of their structures and functions is crucial. Despite significant advances in deciphering proteins’ 3D structures, there remains a pressing need for innovative methodologies to facilitate the computational study of their biochemical and functional properties.
Currently, this challenge is primarily addressed using task-specific models, necessitating extensive data collection and model training for each query. For instance, submissions to the Critical Assessment of Protein Function Annotation algorithms (CAFA) aim to predict the functional annotations (e.g., GO terms) for new protein sequences, as required by the multi-year challenge (Function-SIG, 2024). Similarly, several other task-specific models have been developed to predict specific biochemical properties of proteins, such as ligand binding (Wei et al., 2022) or thermal stability (Blaabjerg et al., 2023).
In this work, we introduce the task of zero-shot Protein Question Answering (PQA) as a novel endeavor to unify protein sequence-related enquiries. To this end, we have developed the Pika framework, comprised of a curated and unbiased dataset specifically designed for the PQA domain, equipped with scientific Q/A pairs for instructional training as well as a robust benchmarking strategy to enable effective evaluation of PQA models. Alongside, we also propose multimodal large language models as a potential solution to this task. Circumventing the need for custom-designed models or datasets for each distinct task, this approach allows for unconstrained free-form scientific enquiry and open-ended answer generation by the large language model (LLM). This is particularly timely, given the significant strides in natural language processing that have led to the development of powerful LLMs. These models, besides their capability in processing natural questions, also encapsulate a large body of knowledge which could provide further context for the model to provide accurate and coherent responses. This paradigm shift enables a more flexible and efficient exploration of protein functionalities, bypassing the need for extensive model training and data collection for individual queries.
In particular, we train and evaluate two multimodal protein-text architectures combining the ESM2 (Lin et al., 2022) protein language model (PLM) with the Phi-2 LLM (Microsoft, 2023), showcasing the seamless integration of protein sequence analysis with natural language processing. This approach yields accurate responses to scientific enquiries regarding the biochemical and functional properties of previously unseen proteins based purely on their sequences, representing an important step forward in this domain. It is also essential to highlight that, our research, leveraging the robust yet modestly scaled Phi-2 LLM, serves as a robust proof-of-concept to demonstrate the potential of this approach. Our results suggest that the strategic adoption of this methodology, especially with more advanced LLMs, has the potential to challenge the current state-of-the-art in task-specific models. Consequently, our approach, by facilitating scientific enquiries through natural language question answering, sets the stage for broader accessibility and efficiency in future research endeavors. This development promises to expedite the discovery process and hypothesis generation, thus reshaping the landscape of biological and biomedical research.
2 Related Work
2.1 Multimodal Language Models
Recent advancements in Multimodal Language Models have significantly enriched the interaction between text and other modalities. Central to this progress are models like SimVLM (Wang et al., 2021), Frozen (Tsimpoukelli et al., 2021), Flamingo (Alayrac et al., 2022), PaLI (Chen et al., 2022) for visual content, alongside innovations like AudioLM (Borsos et al., 2023) for audio data. These models blend the strengths of vision encoders built by convolutional neural networks (LeCun et al., 2015) or transformers (Vaswani et al., 2017; Dosovitskiy et al., 2020), with language models to overcome the limitations of traditional single-modality models. Furthermore, the scope of multimodal models can extend beyond vision, audio, and video to include agent actions (Reed et al., 2022), biomedical data (Zhang et al., 2023; Tu et al., 2023) or molecules (Edwards et al., 2022).
2.2 Protein Language Models
This work draws upon a large body of previous research in creating rich representations of proteins for 3D structure prediction or functional classification (ESM family (Rives et al., 2021; Lin et al., 2022) and AlphaFold (Jumper et al., 2021)). The ESM2 family of PLMs have been trained on a large dataset of protein sequences in a self-supervised manner to yield rich representations for protein sequences through learning the “language of proteins”. These models have demonstrated strong performance in various downstream tasks, including protein folding, functional prediction, and disease-variant effect prediction, highlighting their rich and representative embedding space.
The previous work on integrating text with protein sequences has primarily focused on improving the protein representation by leveraging a large body of unstructured biomedical text or knowledge graphs (ProtST (Xu et al., 2023), OntoProtein (Zhang et al., 2022)), with a focus on functional classification or protein sequence generation from textual descriptions. Conversely, ProteinChat (Guo et al., 2023) has utilized LLM as a knowledge retrieval platform to facilitate discussions about 3D structure databases. These efforts have also led to the creation of two datasets, ProtDescribe (Xu et al., 2023) and PDB-QA (Wang et al., 2023). While PQA has been explored in previous and concurrent research, these have primarily been limited to evaluating knowledge retrieval due to the limitations of available datasets and evaluation metrics (details in Section 3.1). As a result, the task of conducting free-form zero-shot scientific enquiry of new protein sequences remains unexplored.
3 PQA: Dataset & Benchmark
In this section, we detail our dataset and benchmark designs for the Pika framework. Starting with identifying limitations of existing protein-text databases, we then outline the essential criteria for a dedicated PQA dataset, followed by a description of the creation process and the quality control of our specialized dataset, Pika-DS. Finally, we introduce several biochemically motivated benchmarks, essential for assessing the scientific accuracy of PQA models.
3.1 Specialized PQA Dataset
The scientific PQA task is aimed at delivering accurate responses to free-form questions based on an unseen protein sequence. This task emerges from the need for scientific exploration of protein functions via natural language question answering. Training and evaluation of multimodal models for the PQA task necessitates comprehensive datasets with scientific textual annotations linked to corresponding protein sequences. Currently two PQA-relevant datasets are available. While these datasets are suitable for their intended tasks, we found them unfit for the specific purpose of scientific PQA for reasons described below:
  1. 1.
    ProtDescribe dataset: This dataset (Xu et al., 2023) contains textual annotations from three fields of uncurated SwissProt entries, risking bias and data leakage due to high sequence similarity and over-representation of certain protein families. Its use of fields that are not directly inferable from protein sequence (e.g., cellular interaction) limits its utility for PQA. Conversely, it misses out on features like catalytic reactions which are often subject of scientific queries. Crucially, lacking Q/As, it fails to directly support PQA assessments. Therefore, while valuable for enhancing protein representations and text-to-protein generation, as intended by the authors, this dataset exhibits significant limitations for the scientific PQA task.
  2. 2.
    PDB-QA dataset: (Guo et al., 2023) With 30 predefined Q/As about 3D structures from the PDB database, while suitable for knowledge retrieval tasks, PDB-QA’s focus on entry-specific details (e.g., submission date, analysis software) renders it ineffective in the context of scientific PQA.
3.2 Dataset Creation
To address the gaps discussed in the previous subsection, we created Pika-DS, the first specialized and unbiased PQA dataset, accompanied with respective biologically relevant benchmarks for model training and evaluation. We designed this dataset to fulfill three key criteria essential for the training and benchmarking of PQA models:
  • Offers an unbiased representation of known protein sequences, mitigating frequency biases prevalent in existing databases.
  • The dataset ensures that the information associated with each protein sequence is expertly curated, allowing for inference solely based on the protein sequence.
  • It includes pertinent benchmarking metrics to assess model performance especially in zero-shot settings.
# total # unique # words # tokens
Proteins 257 K 252 K - 105.3 M
Info fields 1.17 M 319 K 10.8 M 29.5 M
Summaries 1.33 M 861 K 22.5 M 36.4 M
Questions 1.71 M 98 K 15.7 M 14.9 M
Answers 1.71 M 783 K 18.4 M 31.8 M
Table 1: Pika-DS size and token counts. Total and Unique values for Summaries represent the number of sentences.
Following these criteria, Pika-DS was prepared in three phases:
  1. 1.
    We utilized the SwissProt database as a foundational resource. SwissProt, the reviewed section of UniProt, features approximately 570,000 detailed annotations of protein sequences spanning all domains of life and viruses. For each entry, we extracted an expert-curated list of scientific information covering a wide range of subjects including evolutionary, biochemical and functional properties (Section A.1). Removing proteins shorter than 30 amino acids, we obtained 3.7 million information fields for over 548,000 protein sequences.
  2. 2.
    UniProt also offers similarity-based clustering of its entries at various sequence similarity thresholds of 100%, 90%, and 50%, known as UniRef clusters. For instance, protein sequences that share a minimum of 50% pairwise sequence similarity will be assigned the same UniRef50 cluster. Recognizing a strong bias towards more commonly studied protein families in SwissProt (Figure B.1), where nearly 50% of the SwissProt sequences belong to the largest 6% of clusters, we limited Pika-DS to a maximum of two most informative sequences per UniRef50 cluster using a custom algorithm (Section A.2). This resulted in a debiased set of over 257,000 protein sequences and 1.17 million information fields.
  3. 3.
    Lastly, we instructed GPT3.5 to process the information fields for each protein entry to create the Pika-DS’s three main components (example: Table B.1):
    • Summary: A summary of each protein’s functional and biochemical properties, based solely on the provided information excluding the protein’s name.
    • QAs: Several diverse Q/A pairs for each information field, formatted for LLM training.
    • Metrics: Single-word answers to a set of predefined scientific questions serving as the ground-truth for our Biochem-Lite benchmarks (Section 3.4).
    The prompts for all instructions were fine-tuned to ensure GPT3.5’s adherence to using only the given data, summarizing complex biological reactions accurately, and avoiding speculative or ambiguous language. Examples in Section A.3.
Figure 2: Characteristics of PQA dataset. (a) Distribution of token counts for all examples in Pika-DS. (b) Frequency of words in each position in each section of the dataset. Long words are abbreviated (do.=does, mol.=molecule, fun.=function, we.=weight, bel.=belong).
The final Pika-DS comprises 257,167 protein sequences, carefully selected from 185,128 UniRef50 clusters. This dataset is enriched with detailed descriptors for each sequence: a summary statement divided into sentences, multiple scientific Q/A pairs, and the ground-truth answers for our predefined benchmarking questions. This dataset encapsulates over 105 million protein sequence tokens, approximately 36.4 million tokens in textual summaries, and nearly 47 million tokens in Q/A pairs (Table 1).
3.3 Dataset Analysis
In this subsection, we analyze various facets of our finalized PQA dataset to validate its quality and suitability for training and evaluating multimodal models for the purpose of open-ended scientific enquiries of protein sequences.
Protein over-representation Analysis:
We used pre-computed protein embeddings from UniProt database and visualized them using UMAP (Uniform Manifold Approximation and Projection for Dimension Reduction(McInnes et al., 2018)), highlighting the entries that belong to the top 100 largest clusters. Visual comparison of distribution of protein embeddings before and after our Uniref50-based filtering confirms a strong reduction of over-represented protein families in Pika-DS while maintaining the rich diversity of the dataset (Figure B.2).
Token Count Analysis:
In creating the Q/A pairs using GPT3.5, we instructed the model to cover a broad spectrum of queries and to produce detailed answers suitable for training other LLMs, anticipating elaborate responses rather than single-word answers. This was confirmed by our analysis of token counts in the summary sentences, questions, and answers of the final dataset. This revealed that while questions are typically shorter (8.5 tokens on average), the summary sentences and answers are longer and more extensive (on average 27 and 16.5 tokens). This presents an adequate number of tokens per example, providing sufficient context for the training of multimodal PQA models (Figure 2a).
Word Frequency Analysis:
To ensure the diversity of content in summary statements as well as Q/A pairs, we examined the frequency of words in the initial five positions of each of these categories. As anticipated, common words such as “the” and “it”, in summaries, and interrogatives like “What” and “How” in questions, dominated the first three positions. Similarly, answers often began with “the” or “this” as well as “yes/no” structures, followed by the word “protein”. However, the remaining positions demonstrated a significant lexical diversity. This observation confirmed that our dataset does not exhibit a substantial bias, offering the necessary diversity for effective model training (Figure 2b).
These analyses underscores the high quality and diversity of the Pika-DS, affirming its suitability for further research in this domain.
Metric ID Question Example MLP task
label Self-Pika’s response
mw MALE
What is the molecular weight of
this protein?
55808
The molecular weight of this protein is
54988 KDa.
log value
exact cofactor What is a cofactor of this protein? Zn(2+) The cofactor of this protein is Zn(2+). -
is_enzyme F1
Can this sequence be considered
an enzyme?
True Yes classification
location F1 Where is this protein located? Membrane
The sub-cellular location of this protein
is the cell membrane.
classification
binary Q
mean F1
Is this a membrane protein?
Is this a nuclear protein?
Is this a mitochondrial protein?
True
False
False
Yes, it is a single-pass membrane protein.
No, it is localized to the cytoplasm.
No
Table 2: Questions and example responses for Biochem-Lite (Section A.4 provides further details on motivation and score computation for each metric). Self-Pika is our best performing model. Underlined values indicate the extracted entities for score calculation.
3.4 Benchmarks
In this section, we elaborate on our biochemically-focused benchmarking methodologies tailored for evaluating the scientific accuracy of multimodal PQA models.
Motivation & Design Criteria:
Conventional linguistic metrics like BLEU and ROUGE, while useful in general linguistic contexts, often fall short in assessing scientific correctness and show poor correlation with human judgment (Mathur et al., 2020). As a result these metrics are inadequate for assessing the performance of multimododal PQA models. Therefore, going beyond standard linguistic evaluations, we designed a purpose-built benchmarking approach incorporating a set of predefined, biochemically-significant questions, that are specifically selected to:
  1. 1.
    Reflect the biochemical properties of proteins, ensuring that a model’s accuracy in these responses is indicative of its effectiveness in broader scientific enquiries.
  2. 2.
    Span a spectrum of complexity, from straightforward information extraction with minimal linguistic intricacy to advanced linguistic reasoning for identifying pertinent information.
Two-Tiered Benchmarking System:
Alongside the benchmarking questions for the PQA task, we also require a robust strategy to assess the correctness of open-ended answers to them. Three possible assessment strategies could be imagined: (1) Exact matching of statements, (2) Relevant entity extraction, and (3) Comparison using stronger LLMs (xLLMs). While exact statement matching is useful in some research domains, it lacks the rigour required for assessing scientific accuracy, as such we selected entity extraction and xLLM comparison. While xLLM comparison, although ideal, is computationally prohibitive, entity extraction pragmatically balances scientific accuracy with computational efficiency. This balance of performance vs efficiency motivated us to design a two-tiered benchmarking system comprised of a light-weight benchmark employing entity extraction for efficiency, and a rigorous xLLM-based benchmark for scientific fidelity.
Metric ID Question Ground truth Self-Pika’s response
Reaction
What chemical reaction is
catalyzed by this protein?
EC = 5.1.1.1,
L-alanine $\rightarrow$ D-alanine
This protein catalyzes the conversion of
L-alanine to D-alanine.
Domains
What are the functional
domains of this protein?
PLP-binding barrel, Alanine racemase,
Alanine racemase C-terminal domain-like,
pyridoxal phosphate binding
The functional domains of this protein
include the alanine racemase domain,
Alanine racemase C-terminal domain-like,
Alanine racemase, and PLP-binding barrel.
Cofactors
What are the cofactors
of this protein?
pyridoxal 5’-phosphate
The cofactors of this protein are
pyridoxal 5’-phosphate and magnesium ions.
Table 3: Questions and example responses for Biochem-ReAct. Self-Pika is our best performing model.
  1. 1.
    Biochem-Lite: Pika’s light-weight benchmarking involves a set of pre-defined, scientifically relevant questions with simple answers extracted for each protein using GPT3.5 during Pika-DS creation. These questions are designed to cover a range of biochemical and functional properties (e.g. binding to cofactros or cellular localization) while evaluating multimodal PQA performance at various levels of (1) information extraction from protein embeddings, (2) cross-modal information processing by the LLM, or (3) the LLM’s ability to generate relevant responses (Table 2 and Section A.4). Although the ground truth answers to these questions are single-worded, for evaluation, questions are presented in free-form, and the model provides an open-ended response. This response is then processed for rule-based entity extraction and scoring using adequate metrics.
    • For mw MALE, the mean-absolute log error (MALE) of the predicted molecular weight (MW) is:
      $MALE=\frac{1}{N}\sum_{i=1}^{N}|\log_{10}(\hat{MW}_{i})/(MW_{i})|$ (1)
      where $\hat{MW}_{i}$ is the predicted MW, $MW_{i}$ is the ground truth MW, and $N$ is number of examples.
    • For exact cofactor, the score is computed as:
      $\text{Score}_{\text{exact}}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{1}(\exists w\in R_{i}:w\in GT_{i})$ (2)
      where $R_{i}$ is the set of words in response $i$, $GT_{i}$ is the set of ground truth cofactor words for protein $i$, and $\mathbb{1}$ is the indicator function that is 1 for exactly one match and 0 otherwise.
    • For location F1, the F1 score is defined as:
      $F1=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (3)
      The predicted class assignments for the calculation of Precision and Recall are based on the exclusive presence of correct labels (membrane, nucleus, or mitochondrion) in the generated response. Responses lacking or containing multiple labels receive the none class label.
    • For is_enzyme F1, the F1 score is used as defined in Equation 3. Responses are classified as True or False based on the exclusive presence of yes or no, respectively. Any deviation is labeled as none.
    • For binary Q mean F1, we have:
      $F1_{\text{binary Q mean}}=\frac{1}{3}\sum_{q=1}^{3}F1_{q}$ (4)
      where $F1_{q}$ is the F1 score for each binary question with class assignments similar to the is_enzyme F1.
  2. 2.
    Biochem-ReAct: While light-weight benchmarking questions offer a general overview of model’s performance in various scientific aspects, they do not represent real-world scenarios where the scientific fidelity of open-ended responses is vital. Therefore, we also selected three biochemical questions surrounding reaction & activity of the proteins that can be subjects of scientific enquiry into novel protein sequences (Table 3). The complexity of the responses to these questions necessitates assessment using advanced LLMs such as GPT3.5 and GPT4, therefore imposing a significant computational cost. As a result, Biochem-ReAct is mainly suitable for end-point benchmarking of the best-performing models.
Overall, our two-tiered benchmarking system establishes a balanced and robust framework for evaluating multimodal PQA models, ensuring both scientific rigor and computational feasibility.
4 Baselines & Methodology
In this section we outline the baselines, architecture and the training strategy for our proposed solutions to the task of scientific PQA. All baseline models and architectures were implemented using PyTorch Lightning, and the complete codebase is accessible on github.com/EMCarrami/Pika.
4.1 Baselines
The prior research in the subject of multimodal text-protein models has primarily been focused on improving protein embeddings with textual data or for text-guided protein sequence generation (Xu et al., 2023; Zhang et al., 2022). The use of PQA in other studies (Guo et al., 2023) has been limited to knowledge retrieval, which do not provide comparable benchmarking opportunities for scientific PQA, mainly due to the high presence of non-scientific content in the PDB-QA dataset (e.g., author names). Therefore, in the absence of relevant state-of-the-art, here we define four baselines specifically for scientific PQA in the context of Biochem-Lite benchmarking questions (Section 3.1).
  • Random Baseline: Selects a random answer for each question from the pool of all possible answers to the same question within the evaluated dataset/split.
  • MLP: An MLP on protein sequence embeddings of the PLM (Section A.5). Applies to classification or value prediction tasks in Biochem-Lite (Table 2). It sets an upper limit for the information content of PLM embeddings, with any improvements potentially indicating generalization due to the knowledge encapsulated in the LLM.
  • LLM Only: Pre-trained LLM’s base response to Biochem-Lite questions. Phi-2 was used.
  • Pika w/o PLM: Mirrors Self-Pika’s architecture and training, substituting the PLM with simple token embeddings of the protein sequences. It defines the base performance for the Biochem-Lite benchmarks (except for “mw MALE”).
Figure 3: Schematic representation of Cross- and Self-Pika architectures for the scientific PQA task. PLM = Protein Language Model (protein sequence encoder), LLM = Large Language Model. Only the adapter and cross-attention modules (both in green) are trained.
Models Biochem-Lite Biochem-ReAct
mw
MALE
exact
cofactor
is_enzyme
F1
location
F1
binary
mF1
pplx Reaction Domains Cofactor
Random 0.35 0.13 0.49 0.34 0.46 - - - -
Baselines MLP 0.07 - 0.86 0.89 - - - - -
Phi-2 only 4.53 0.03 0.34 0.00 0.31 - - - -
Pika w/o PLM 0.02 0.21 0.7 0.47 0.21 2.30 0.01 0.32 0.82
Cross-Pika (Q) 0.02 0.47 0.89 0.77 0.13 1.69 0.36 0.60 0.94
Our Models Self-Pika (Q) 0.04 0.54 0.89 0.76 0.44 1.98 0.58 0.79 0.97
Self-Pika (Q+S) 0.04 0.57 0.88 0.76 0.57 2.23 0.53 0.75 0.96
Table 4: Performance of Pika models on Biochem-Lite and Biochem-ReAct metrics. Bold values indicate the best score. Standard deviation for all scores $<0.05$ except for those indicated by underline. Biochem-Lite results on val set and from three different seeded training. Biochem-ReAct results on test set and a single seed training. “binary mF1” = “binary Q mean F1”, pplx = perplexity, Q = training with QA data only, Q+S = training with QA and Summary data.
4.2 Models
We propose two robust multimodal architectures for PQA tasks, drawing inspiration from successful VLM strategies like Frozen (Tsimpoukelli et al., 2021), Flamingo (Tsimpoukelli et al., 2021) and Prismer (Liu et al., 2023). In both architectures we keep the pre-trained LLM and PLM frozen, only incorporating trainable “adapter” modules that extract relevant latent representations from PLM’s output (protein sequence embeddings) and integrate them into the LLM’s transformer architecture, facilitating seamless transfer of information across modalities.
  • Cross-Pika uses multiple independent learnable adapters, each creating distinct protein latent embeddings for each transformer layer of the LLM. These embeddings are injected into the LLM using a gated cross-attention mechanism before each native self-attention layer (Fig. 3). At the cost of increased complexity, this design allows for a nuanced and layer-specific modulation of the LLM, potentially enabling it to process complex protein-related information more effectively, which could aid the model in answering biochemically intricate PQA queries.
  • Self-Pika uses a single learnable adapter to transform protein sequence embeddings into latent embeddings compatible with language token embeddings. The transformed embeddings, concatenated at the beginning of the LLM’s initial token embeddings, allow the protein latent embeddings to influence the LLM’s response through its internal self-attention mechanisms. This approach effectively creates soft-prompts conditioned on the input protein sequence (Fig. 3). This architecture simplifies the integration process, reducing computational overhead while allowing for the incorporation of essential protein characteristics into the LLM’s response.
Considering the diversity in protein lengths, we opted for the Perceiver architecture (Jaegle et al., 2021) as the learnable adapter for both models. This choice standardizes the transformation of protein embeddings into a consistent number of latent embeddings for seamless cross-modality information transfer. Optionally, predefined placeholder tokens can also be added to each question to signaling the LLM about the presence of the protein sequence. In this work we used ESM2 (Lin et al., 2022) as the pre-trained PLM and Phi-2 (Microsoft, 2023) or GPT2 (Radford et al., 2019) as the pre-trained LLM. Our best performing model is based on the Self-Pika architecture, with ESM2-L (esm2_33t_650M) and Phi-2 pre-trained models.
4.3 Training
We devised a simple training strategy for PQA models, keeping both the PLM and LLM frozen and optimizing for the causal language model loss with AdamW. Hyperparameters were optimised following a greedy search as detaile in section A.6.
All training was performed on a single A100-80GB or H100-80GB Nvidia GPU. Unless otherwise stated, the QA section of the Pika-DS was used for the training of all Pika models. Examples were split to train, validation and test set based on the UniRef50 cluster of their respective protein sequences. As a result, protein sequence from the evaluation sets (validation and test) are never present in training data.
Training Mode Biochem-Lite Biochem-ReAct
Model S Ctrl Q QA
mw
MALE
exact
cofactor
binary Q
mean F1
Reaction Domains Cofactor
self-Pika (S) 4.54 0.01 0.00 - - -
self-Pika (S+C) 2.30 0.32 0.42 0.22 0.70 0.95
self-Pika (Q) 0.04 0.54 0.44 0.58 0.79 0.97
self-Pika (Q+S) 0.04 0.57 0.57 0.53 0.75 0.96
Table 5: Perfomance of Self-Pika in the absence of Q/As during training. Bold values indicate best score. Standard deviation for all scores $<0.05$ except for those indicated by underline. S = Summary, Ctrl Q = Control Question, QA = Q/A pairs.
5 Experiments & Results
We start this section by validating our benchmarking strategy, followed by the zero-shot evaluation of best models. In the end, we present the results of our ablation study, highlighting the key factors in scientific PQA model performance.
5.1 Benchmarking Results
Biochem-Lite Results:
Table 4 summarizes the benchmarking results of Pika models in comparison with various baselines. It is evident that both Self- and Cross-Pika architectures outperform all lower-bound baselines, indicating successful cross-modality information transfer to the LLM. Notably, our top models also surpass the upper-bound MLP baseline on key Biochem-Lite benchmarking questions, underscoring the potential for generalization through the LLM’s knowledge. Additionally, we assessed the reliability of ROUGE scores for evaluating PQA model performance. ROUGE scores remained consistent, regardless of whether Self-Pika utilized simple protein token embeddings (w/o PLM) or embedded protein sequences (Figure B.3). This finding confirms the inadequacy of traditional linguistic metrics for the scientific PQA task, thus demonstrating the utility of our Biochem-Lite metrics.
Biochem-ReAct Results:
Our Biochem-ReAct benchmarking, aimed at measuring scientific fidelity, yielded promising results (Table 4). The Self-Pika model demonstrated strong performance across all three Biochem-ReAct questions, exceeding the without PLM baseline by 57%, 31%, and 15% for correctly identifying Reactions, Domains, and Cofactors, respectively, for previously unseen protein sequences. Additionally, comparing Biochem-Lite results against Biochem-ReAct scores for 12 high-performing checkpoints with varied configurations revealed that while traditional linguistic metrics, including perplexity, fall short in predicting the real-world efficacy of PQA models, two Biochem-Lite questions—“exact cofactor recall” and “binary Q mean F1”—emerged as strong indicators of multimodal PQA models’ real-world performance (Figure B.4).
Overall, these results validate our benchmarking strategy and establish Self-Pika as the initial state-of-the-art for the task of scientific PQA.
5.2 Zero-shot PQA:
We define zero-shot PQA as answering novel questions about unseen protein sequences. Given the unbiased nature of Pika-DS and data splits based on UniRef50 IDs, our validation and test sets largely comprise unseen protein sequences. Moreover, the “binary Q” component in Biochem-Lite consists of questions never included in the Pika-DS, suggesting our benchmarking broadly represents zero-shot PQA tasks. This section delves into stricter zero-shot scenarios to better understand PQA model behavior.
PQA Learning without Questions:
To understand the significance of Q/A pairs in the training set, we conducted training under two stringent conditions. The first involved using only summary sentences as labels for proteins without any Q/As. The second condition included summary sentences and a single Control Question, “Is this a real protein?”. During training, we randomly shuffled the tokens within the sequence of half the proteins, setting the expected answer to this question as No (for summary labels all protein sequences remained unchanged). This aims to train the model to understand the task of question answering, with summaries providing the scientific context for learning. Remarkably, introducing the Control Question, even when no other questions were present during training, significantly enhanced performance (Table 5). Additionally, we observed that the performance of the model on the Control Question follows a similar pattern as other metrics (Table 6). These observation suggest that the current bottleneck in PQA performance is likely the LLM and extending this work to larger LLMs could further improve performance and generalization capabilities.
Random
LLM
only
self-Pika
w/o PLM
self-Pika
F1 Score 0.49 0.00 0.43 0.99
Accuracy 0.50 0.00 0.51 0.99
Table 6: Performance on Control Question.
LLM PLM
mw
MALE
exact
cofactor
is_enzyme
F1
location
F1
binary Q
mean F1
perplexity
ESM2-S 0.14 0.22 0.84 0.65 0.26 3.26
GPT2-M ESM2-M 0.06 0.26 0.86 0.64 0.33 3.12
ESM2-L 0.07 0.32 0.86 0.70 0.23 3.05
ESM2-S 0.02 0.51 0.86 0.71 0.33 1.99
Phi-2 ESM2-M 0.02 0.56 0.88 0.75 0.46 1.95
ESM2-L 0.04 0.54 0.89 0.76 0.44 1.98
Table 7: Effect of LLM and PLM size on Self-Pika models’ performance. Standard deviation for all scores $<0.05$ except for those indicated by underline.
PQA Learning with Novel Proteins:
In training multimodal Pika models, we ensured that proteins similar to those in the training dataset were excluded from the test and validation sets, effectively creating a zero-shot scenario for our benchmarking. Given that the ESM2 model was pre-trained on a vast dataset of proteins, there was a potential concern that its embeddings might be influenced by previously encountered sequences. To mitigate this, we utilized the evaluation split reported for ESM2’s pre-training, categorizing our benchmarking results into proteins seen and unseen by ESM2. The analysis revealed no significant performance difference between these conditions, affirming that Pika’s zero-shot capabilities are not compromised by the prior knowledge embedded in ESM2. This finding underscores the robustness of Pika models in genuine zero-shot PQA tasks, independent of ESM2’s pretraining exposure.
5.3 Ablation Studies
To identify key contributing factors, we compared model performance across various dimensions. Most notably, reducing the size of the LLM from Phi-2, which has 2.8 billion parameters, to GPT-2 Medium, with 355 million parameters, resulted in a 20% decrease in the key “exact cofactor” metric (Table 7). However, the size of the ESM2 model did not seem to have a significant impact, with performance remaining largely similar between ESM2-S, ESM2-M, and ESM2-L (8M, 65M and 350M parameters, respectively). This observation could be attributed to the richness of representations in the Pre-trained PLMs, potentially highlighting a bottleneck in the LLM or the quality and quantity of the data (Table 7). Concordantly, the comparison of results from the zero-shot experiment also confirms that with a lower volume of data, a tailor-made dataset could yield better results. However, as the dataset size increases—a scenario analogous to the image domain—this factor becomes less crucial. These findings suggest that while LLM capacity significantly influences PQA model performance, the sophistication of sequence embeddings provided by ESMs reaches a point of diminishing returns, underscoring the importance of focusing on LLM enhancements and data quality for future improvements.
6 Conclusion
We’ve established a pioneering framework for zero-shot PQA, presenting a significant step forward in the application of LLMs for scientific enquiry. The introduction of specialized datasets and biologically relevant benchmarks underpins future explorations at the intersection of computational biology and artificial intelligence. Key insights from our zero-shot evaluations and ablation studies highlight the crucial role of LLMs in multimodal PQA performance, underscoring the importance of the knowledge depth and linguistic capabilities of LLMs. In essence, this work not only sets a new standard for scientific PQA but also opens the door for the research community to further refine and expand upon our methodologies. Looking ahead, leveraging larger, more diverse LLMs may offer substantial gains in model generalization, driving us towards the goal of automating accurate scientific enquiry.
7 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning and Biology. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgements
We would like to thank Dr Aida Nematzadeh for critical advice and Maryam Karaminejad Ranjbar for graphic design.
Appendix A Supplementary Methods
A.1 Data collection from UniProt
We retrieved XML files of all SwissProt entries longer than 30 amino acids, with the cut-off date of 14-08-2023, using UniProt’s API. Since these entries contained extensive extraneous information, like author names, submission dates, etc., we employed a rule-based pre-processing approach to extract fields relevant to each protein’s functional and biochemical characteristics. This selection, guided by an expert biochemist, included sequence (molecular mass and length), organism (top three taxonomic levels), catalytic activity (including EC number), biophysicochemical properties (pH and temperature dependence), cofactor, subunit (excluding fields containing “interact”, “associate”, or “complex”), subcellular location (excluding isoforms), and functional domains: GO (only molecular function, omitting biological process and cellular component), Gene3D, and SUPFAM.
A.2 Debiasing SwissProt Entries Using UniRef50 Clusters
To mitigate bias in SwissProt entries, we employed UniRef50 clusters, sourced from the UniProt FTP’s idmapping file. First, we merged clusters representing isoforms with their corresponding main protein clusters to consolidate isoform-driven redundancy. Specifically, if the isoform of a protein was a member of another Uniref50 cluster, we merged the isoform cluster into the main protein’s cluster. This step aimed to tighten clustering criteria, avoiding oversampling due to isoforms. Next, within each merged cluster, sequences shorter than 25% of their cluster’s median length were excluded to ensure a focus on sufficiently representative sequences. This filtering is due the fact that some Uniprot sequences have incomplete sequences. Finally, following the methodology outlined by Jiang et al. (Jiang et al., 2022), we calculated gzip information for each entry. Initially, the entry with the highest gzip information was identified. Subsequently, we assessed the additional gzip information provided by each remaining entry relative to this top candidate. The entry offering the highest additional gzip information was selected for inclusion. Throughout this selection process, cluster representatives were given priority in the event of a tie, ensuring the most informative representatives were chosen. This debiasing process was designed to refine the Pika-DS by leveraging UniRef50 clusters, enhancing the dataset’s quality and representativeness for downstream applications.
A.3 Example Instructions to GPT3.5
You will receive details about a specific protein. Perform the following tasks: 1) Provide a factual summary, with a maximum of 500 words, that accurately and scientifically describes the functional, biochemical and structural properties of this protein based only on the provided information. Ensure that the summary follows a natural and scientific flow, starting with general information such as structure, localization and taxonomy before detailing functional and biochemical properties. Ensure that all key points are covered and DON’T provide any further information than what is stated in the input. 2) For each type of information provided, create a question-and-answer pair to elucidate an aspect of the protein’s functionality or biochemical properties without using the protein’s name.
Use the below ; separated two-element tuples structure for the output: (summary: this protein…); (what is question 1?, the answer 1); (what is question 2?, the answer 2); …
For both tasks if the input contains large group of properties only provide the canonical and crucial information rather than enumerating every single entry. DON’T use any of your knowledge to add additional context or information. DON’T add any speculative or unwarranted information that is not specifically provided. AVOID using generic phrases or embellished terms such as ’highly’, ’several’, ’diverse’ and ’various’.
A.4 Biochem-Lite Metric Questions and Motivation
mw MALE
  • Question: What is the molecular weight of this protein?
  • Score: Mean Absolute Error of log values
  • Presence in training: in exact form
  • Motivation: Proteins are large polymers of amino acids. The molecular weight of a protein depends not only on the number of amino acids but also on the types of amino acids in the sequence.
Cofactor recall
  • Question: What is a cofactor of this protein?
  • Score: If exact match, score=1; else 0 (average across examples)
  • Presence in training: in exact form
  • Motivation: Many proteins require other small molecular entities to function. These could be ions, or small molecules, or a combination of these, such as a Heme-Fe. Proteins often bind to their cofactors via motifs, which are short 3D structures formed due to the presence of specific sequences. Identifying the cofactor of a protein requires the extraction and analysis of these sequence motifs from the sequence embeddings.
is_enzyme F1
  • Question: Can this sequence be considered an enzyme?
  • Score: F1 score
  • Presence in training: in similar form
  • Motivation: The motivation for this question is similar to that of cofactor recall, focusing on the functional characterization of proteins. Enzymatic activity is a crucial aspect of protein function, requiring specific structural or sequence features for identification.
Location F1
  • Question: Where is this protein located?
  • Score: F1 score
  • Presence in training: in exact form
  • Motivation: Proteins must be localized to specific compartments within cells to perform their functions correctly. This localization is directed by specific targeting sequences or 3D structures. Extracting this information can be challenging, and the identification of the destination for a protein can vary depending on the organism and context.
Binary localization average F1
  • Question: Is this a membrane protein? Is this a nuclear protein? Is this a mitochondrial protein?
  • Score: Mean of F1 scores of all 3 questions
  • Presence in training: never in this form
  • Motivation: This question is related to the previous one but focuses on specific locations in a binary format. While the information extraction process is similar, the LLM needs to perform reasoning to find the answer.
A.5 MLP Baseline
We use a 3-layer MLP with GELU activation on protein sequence embeddings from the Pre-trained PLM. Where possible, we convert each light-weight benchmarking question into a classification or value prediction task (Table 2). The MLP is then trained to predict the correct class labels or values using cross entropy or mean squared error loss, respectively. This baseline sets an upper limit for the information content inherent in PLM embeddings, with any improvements potentially indicating generalization due to the knowledge encapsulated in the LLM.
A.6 Hyperparameter Optimization
Greedy hyper parameter search was performed for both architectures, monitoring the is_enzyme metric. For all experiments, training was performed on 25000 protein sequences and metrics were computed on 250 unseen proteins with gpt2-medium model as the LLM and esm2_t12_35M model as the protein sequence encoder. The sweep was performed, in order, on optimizer’s weight decay [0, 1e-4, 1e-2] and learning rate [1e-5, 1e-4, 1e-3], batch size [2, 4, 8], Perceiver latent size [32, 64, 100] and the number of Perceiver layers [1, 2, 4]. The final set of hyper parameters were as follows: learning rate: 1e-4, weight decay: 1e-4, batch size: 8, Perceiver latent size: 100, and number of Perceiver layers: 1 (Cross-Pika architecture), 4 (Self-Pika architecture).
Appendix B Supplementary Figures and Tables
uniref UniRef50_Q9HUN4
sequence MRPLVATVDLTALRHNYLLAKQCAPQRKAFAVVKANAYGHGAPEAVTALREIADGFAVACLEEAE[…]
fields catalytic activity: EC = 5.1.1.1, L-alanine = D-alanine
cofactor: pyridoxal 5’-phosphate
functional domains: alanine racemase activity, Alanine racemase, Alanine racemase C-terminal domain-like,
PLP-binding barrel, pyridoxal phosphate binding
taxonomy: Bacteria, Pseudomonadota, Gammaproteobacteria
protein size: 358 aa, 38485 KDa
summary
This protein is an alanine racemase enzyme found in various bacteria, specifically in the
Pseudomonadota phylum of the Gammaproteobacteria class
It is composed of 358 amino acids with a molecular weight of 38,485 KDa
The protein exhibits a catalytic activity with an Enzyme Commission (EC) number of 5.1.1.1,
where it converts L-alanine to D-alanine
Its enzymatic function is facilitated by a cofactor known as pyridoxal 5’-phosphate
The protein contains several functional domains, including the Alanine racemase activity domain,
the Alanine racemase C-terminal domain-like, the PLP-binding barrel domain, and
the pyridoxal phosphate binding domain.
qa
What is the catalytic activity of this protein?
The catalytic activity of this protein is the conversion of L-alanine to D-alanine.
What is the cofactor associated with this protein?
The cofactor associated with this protein is pyridoxal 5’-phosphate.
Which phylum does this protein belong to?
This protein belongs to the Pseudomonadota phylum.
What is the molecular weight of this protein?
The molecular weight of this protein is 38,485 KDa.
How many amino acids are present in this protein?
This protein is composed of 358 amino acids.
What are the functional domains of this protein?
The functional domains of this protein include the Alanine racemase activity domain,
the Alanine racemase C-terminal domain-like, the PLP-binding barrel domain, and
the pyridoxal phosphate binding domain.’,
What class of bacteria is this protein found in?
This protein is found in the Gammaproteobacteria class.
Can this protein act as an enzyme?
Yes
metrics in_membrane False
in_nucleus False
in_mitochondria False
is_enzyme True
cofactor pyridoxal 5’-phosphate
Table B.1: An example entry in Pika-DS representing Uniprto ID A4VQM5
Figure B.1: Over-representation bias in SwissProt database.
Figure B.2: Comparison of sequence bias in SwissProt database, before and after filtering. Members of the top100 largest UniRef50 clusters are colored. The strong overrepresentation of highly studied protein groups is apparent in the unfiltered plot, while the Pika-DS shows a significant reduction in this bias
Figure B.3: Evaluating the Effectiveness of Biochem-Lite Versus Traditional Linguistic Metrics for Scientific Accuracy of PQA. Statistical significance is determined through a one-tailed paired t-test across three randomly seeded data subsets and model training (significance guide: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$). By design, Self-Pika is expected to outperform Pika w/o PLM across all metrics, except for the mw MALE metric.
Figure B.4: Cross-correlation of Biochem-ReAct Benchmarks with Biochem-Lite/Linguistic Benchmarks. Absolute values.

### v2

@@
-PQA: Zero-shot Protein Question Answering for Free-form Scientific Enquiry with Large Language Models
+PQA: Zero-shot Protein Question Answering for Free-form Scientific Enquiry with Language Models
 Abstract
-We introduce the novel task of zero-shot Protein Question Answering (PQA) for free-form scientific enquiry.
-Given a previously unseen protein sequence and a natural language question, the task is to deliver a scientifically accurate answer.
-This task not only supports future biological research, but could also provide a test bed for assessing the scientific precision of large language models (LLMs).
-We contribute the first specialized dataset for PQA model training, containing 257K protein sequences annotated with 1.97M scientific question-answer pairs.
-Additionally, we propose and study several novel biologically relevant benchmarks for scientific PQA.
-Employing two robust multi-modal architectures, we establish an initial state-of-the-art performance for PQA and reveal key performance factors through ablation studies.
-Our comprehensive PQA framework, named Pika, including dataset, code, model checkpoints, and a user-friendly demo, is openly accessible on github.com/EMCarrami/Pika, promoting wider research and application in the field.
-Machine Learning, Multimodel Large Language Models, Protein Sequence
-11footnotetext: Relation Therapeutics, London, UK22footnotetext: Ludwig Maximilian University of Munich, Germany*Corresponding author: Eli M Carrami, eli@relationrx.com, eli.carrami@gmail.com
+Understanding protein structure and function is crucial in biology.
+However, current computational methods are often task-specific and resource-intensive.
+To address this, we propose zero-shot Protein Question Answering (PQA), a task designed to answer a wide range of protein-related queries without task-specific training.
+The success of PQA hinges on high-quality datasets and robust evaluation strategies, both of which are lacking in current research.
+Existing datasets suffer from biases, noise, and lack of evolutionary context, while current evaluation methods fail to accurately assess model performance.
+We introduce the Pika framework to overcome these limitations.
+Pika comprises a curated, debiased dataset tailored for PQA and a biochemically relevant benchmarking strategy.
+We also propose multimodal large language models as a strong baseline for PQA, leveraging their natural language processing and knowledge.
+This approach promises a more flexible and efficient way to explore protein properties, advancing protein research.
+Our comprehensive PQA framework, Pika, including dataset, code, and model checkpoints, is openly accessible on Github, promoting wider research in the field.
 1 Introduction
-Figure 1: Schematic of Pika framework. Pika-DS is created from filtered SwissProt entries followed by processing using GPT3.5.
 Proteins, essential to biological functions, are complex macromolecules that perform a myriad of cellular roles determined by their complex structures and interactions.
 As a polymer of amino acids, drawn from a pool of 20 natural ones, the sequence in which these building blocks are arranged dictates the protein’s three-dimensional structure, which is critical for its function.
-These structures facilitate interactions with various molecules, underpinning numerous biological processes.
 Given the critical role of proteins in both fundamental biology and applied biomedical research, a deep understanding of their structures and functions is crucial.
 Despite significant advances in deciphering proteins’ 3D structures, there remains a pressing need for innovative methodologies to facilitate the computational study of their biochemical and functional properties.
-Currently, this challenge is primarily addressed using task-specific models, necessitating extensive data collection and model training for each query.
-For instance, submissions to the Critical Assessment of Protein Function Annotation algorithms (CAFA) aim to predict the functional annotations (e.g., GO terms) for new protein sequences, as required by the multi-year challenge (Function-SIG, 2024).
+Currently, training individual models tailored to specific tasks are the primary approach to computationally studying the biochemical and functional properties of proteins, requiring extensive data collection and training for each unique task.
+For instance, submissions to the Critical Assessment of Protein Function Annotation algorithms (CAFA) aim to predict the functional annotations such as GO terms for new protein sequences, as required by the multi-year challenge (Function-SIG, 2024).
 Similarly, several other task-specific models have been developed to predict specific biochemical properties of proteins, such as ligand binding (Wei et al., 2022) or thermal stability (Blaabjerg et al., 2023).
-In this work, we introduce the task of zero-shot Protein Question Answering (PQA) as a novel endeavor to unify protein sequence-related enquiries.
-To this end, we have developed the Pika framework, comprised of a curated and unbiased dataset specifically designed for the PQA domain, equipped with scientific Q/A pairs for instructional training as well as a robust benchmarking strategy to enable effective evaluation of PQA models.
-Alongside, we also propose multimodal large language models as a potential solution to this task.
-Circumventing the need for custom-designed models or datasets for each distinct task, this approach allows for unconstrained free-form scientific enquiry and open-ended answer generation by the large language model (LLM).
-This is particularly timely, given the significant strides in natural language processing that have led to the development of powerful LLMs.
-These models, besides their capability in processing natural questions, also encapsulate a large body of knowledge which could provide further context for the model to provide accurate and coherent responses.
-This paradigm shift enables a more flexible and efficient exploration of protein functionalities, bypassing the need for extensive model training and data collection for individual queries.
-In particular, we train and evaluate two multimodal protein-text architectures combining the ESM2 (Lin et al., 2022) protein language model (PLM) with the Phi-2 LLM (Microsoft, 2023), showcasing the seamless integration of protein sequence analysis with natural language processing.
-This approach yields accurate responses to scientific enquiries regarding the biochemical and functional properties of previously unseen proteins based purely on their sequences, representing an important step forward in this domain.
-It is also essential to highlight that, our research, leveraging the robust yet modestly scaled Phi-2 LLM, serves as a robust proof-of-concept to demonstrate the potential of this approach.
+To address the limitations of current approaches, we propose to unify protein sequence-related enquiries under a more generic task of zero-shot Protein Question Answering (PQA), where there can be free-form inquiries about known or novel protein sequences.
+The development and assessment of effective zero-shot PQA models critically depend on the availability of high-quality datasets and robust evaluation strategies, both of which present significant challenges in the current research landscape; unfortunately, existing datasets available for related tasks often exhibit significant limitations.
+They frequently demonstrate biases towards specific protein families, overlook the critical evolutionary relationships between proteins, and may contain noisy or unreliable annotations.
+These shortcomings hinder the development of effective PQA models that can generalize across diverse proteins and accurately answer complex queries.
+Similarly, the robust evaluation of PQA models necessitates biochemically-relevant benchmarking strategies that focus on the scientific accuracy of model predictions within the context of the specific questions posed.
+Previous research efforts have not adequately addressed these challenges.
+They often neglect the impact of evolutionary relationships on information leakage, which can lead to overestimation of model performance.
+Moreover, they predominantly rely on metrics like BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004), which have been consistently shown to be inadequate for assessing the accuracy of scientific statements (Mathur et al., 2020).
+This lack of suitable evaluation tools further impedes progress in the field.
+To address these limitations, we have developed the Pika framework, comprised of a curated and debiased dataset specifically designed for the PQA domain, equipped with scientific question and answer (QA) pairs for instructional training as well as a robust and biochemically relevant benchmarking strategy to enable effective evaluation of PQA models.
+Alongside, we also propose multimodal large language models (LLMs) as a potential solution to PQA and create evolution-aware splits for assessment of their performance.
+Besides their capability in processing natural questions, LLMs encapsulate a large body of knowledge which could provide further context for the model, enabling it to perform a more flexible and efficient exploration of protein functionalities, bypassing the need for extensive model training and data collection for individual queries.
+Finally, comparing the performance of multimodal LLMs with various relevant baselines on our evolution-aware data splits, we show that these models are a promising direction for PQA.
+In particular, we train and evaluate two multimodal protein-text architectures combining the ESM2 (Lin et al., 2022) protein language model (PLM) with the Phi-2 LLM (Microsoft, 2023), showcasing the seamless integration of protein sequence analysis with natural language processing while highlight the shortcomings of these models in particular when dealing with evolutionary distant proteins.
 Our results suggest that the strategic adoption of this methodology, especially with more advanced LLMs, has the potential to challenge the current state-of-the-art in task-specific models.
-Consequently, our approach, by facilitating scientific enquiries through natural language question answering, sets the stage for broader accessibility and efficiency in future research endeavors.
-This development promises to expedite the discovery process and hypothesis generation, thus reshaping the landscape of biological and biomedical research.
+It is essential to highlight that, our research, leveraging the robust yet modestly scaled Phi-2 LLM, serves as a robust proof-of-concept to demonstrate the potential of this approach.
+We hope our work fuels further research and opens new avenues towards performant PQA models.
 2 Related Work
-2.1 Multimodal Language Models
-Recent advancements in Multimodal Language Models have significantly enriched the interaction between text and other modalities.
-Central to this progress are models like SimVLM (Wang et al., 2021), Frozen (Tsimpoukelli et al., 2021), Flamingo (Alayrac et al., 2022), PaLI (Chen et al., 2022) for visual content, alongside innovations like AudioLM (Borsos et al., 2023) for audio data.
-These models blend the strengths of vision encoders built by convolutional neural networks (LeCun et al., 2015) or transformers (Vaswani et al., 2017; Dosovitskiy et al., 2020), with language models to overcome the limitations of traditional single-modality models.
-Furthermore, the scope of multimodal models can extend beyond vision, audio, and video to include agent actions (Reed et al., 2022), biomedical data (Zhang et al., 2023; Tu et al., 2023) or molecules (Edwards et al., 2022).
-2.2 Protein Language Models
-This work draws upon a large body of previous research in creating rich representations of proteins for 3D structure prediction or functional classification (ESM family (Rives et al., 2021; Lin et al., 2022) and AlphaFold (Jumper et al., 2021)).
-The ESM2 family of PLMs have been trained on a large dataset of protein sequences in a self-supervised manner to yield rich representations for protein sequences through learning the “language of proteins”.
-These models have demonstrated strong performance in various downstream tasks, including protein folding, functional prediction, and disease-variant effect prediction, highlighting their rich and representative embedding space.
-The previous work on integrating text with protein sequences has primarily focused on improving the protein representation by leveraging a large body of unstructured biomedical text or knowledge graphs (ProtST (Xu et al., 2023), OntoProtein (Zhang et al., 2022)), with a focus on functional classification or protein sequence generation from textual descriptions.
-Conversely, ProteinChat (Guo et al., 2023) has utilized LLM as a knowledge retrieval platform to facilitate discussions about 3D structure databases.
-These efforts have also led to the creation of two datasets, ProtDescribe (Xu et al., 2023) and PDB-QA (Wang et al., 2023).
-While PQA has been explored in previous and concurrent research, these have primarily been limited to evaluating knowledge retrieval due to the limitations of available datasets and evaluation metrics (details in Section 3.1).
-As a result, the task of conducting free-form zero-shot scientific enquiry of new protein sequences remains unexplored.
-3 PQA: Dataset & Benchmark
-In this section, we detail our dataset and benchmark designs for the Pika framework.
-Starting with identifying limitations of existing protein-text databases, we then outline the essential criteria for a dedicated PQA dataset, followed by a description of the creation process and the quality control of our specialized dataset, Pika-DS.
-Finally, we introduce several biochemically motivated benchmarks, essential for assessing the scientific accuracy of PQA models.
-3.1 Specialized PQA Dataset
+Since PQA is distinct from other question-answering tasks, existing scientific QA datasets such as ScienceQA (Lu et al., 2022), PubMedQA (Jin et al., 2019) or SQuAD (Rajpurkar et al., 2016, 2018) are not suitable for PQA training or evaluation.
+This is because unlike standard QA tasks where the answer to the question can be inferred using logic or existing knowledge, in PQA the answer to the question must be extracted from cross-modality embeddings via the LLM, therefore necessitating specialized protein-text datasets that annotate protein sequences with relevant labels.
+Previous studies on integrating text and protein sequences have primarily leveraged large-scale unstructured biomedical text or knowledge graphs to refine protein representations, facilitating downstream tasks like functional classification and sequence generation (e.g.
+ProtST (Xu et al., 2023), OntoProtein (Zhang et al., 2022)).
+Alongside, ProteinChat (Guo et al., 2023) has utilized LLMs for knowledge retrieval to facilitate discussions about 3D structure databases, and concurrent to our work, Mol-Instructions (Fang et al., 2024) focuses on improving LLM’s understanding of biomolecules (e.g., proteins) by fine-tuning existing LLMs.
+These recent efforts while introducing valuable protein-text datasets have limitations that make them unsuitable for scientific PQA, as detailed below:
+  1. 1.
+    ProtDescribe (Xu et al., 2023): This dataset contains textual annotations from three fields of uncurated SwissProt entries, risking bias and data leakage due to high sequence similarity and over-representation of certain protein families (Fig.
+B.1).
+Also, its use of fields that are not directly inferable from protein sequence alone (e.g., cellular interaction) limits its utility for PQA.
+Conversely, it misses out on features like catalytic reactions which are often subject of scientific queries.
+Crucially, lacking QAs, it fails to directly support PQA assessments.
+Therefore, while valuable for enhancing protein representations and text-to-protein generation, as intended by the authors, this dataset is not suitable for the scientific PQA.
+  2. 2.
+    PDB-QA (Guo et al., 2023) : With a limited set of 30 predefined QAs on 3D structures from the PDB database, while suitable for knowledge retrieval tasks, its focus on entry-specific details (e.g., submission date, analysis software) renders it ineffective in the context of scientific PQA.
+  3. 3.
+    Mol-Instructions (Fang et al., 2024): This dataset was released concurrent to our work.
+The protein section of Mol-Instructions represents a relevant set of template-based textual annotations for proteins tailored towards five downstream tasks, including question-answering.
+The dataset is curated from SwissProt and is debiased using a 90% similarity threshold.
+This lenient threshold results in an abundance of highly related proteins, which leads to bias and in the absence of an evolution-aware splitting strategy could cause leakage across data splits.
+Furthermore, in all previous and concurrent research authors have relied on BLEU or ROUGE metrics for evaluating the scientific accuracy of generate captions or responses.
+However, these metrics are not suitable for assessing accuracy (Mathur et al., 2020), therefore limiting the scientific scope of past research in PQA domain.
+As a result, due to the lack of relevant benchmarking strategies as well as biases and noise in existing datasets, the task of free-form zero-shot scientific enquiry of new protein sequences remains unexplored.
+Figure 1: Schematic of Pika framework. Pika-DS is created from filtered SwissProt entries followed by processing using GPT3.5.
+3 Pika Framework
+Here we detail our dataset, benchmarking and baseline designs for the Pika framework (Fig.
+1).
+All baselines, model architectures and benchmarking were implemented in PyTorch Lightning , and the complete codebase is accessible on github.com/EMCarrami/Pika.
+3.1 Pika Dataset
 The scientific PQA task is aimed at delivering accurate responses to free-form questions based on an unseen protein sequence.
 This task emerges from the need for scientific exploration of protein functions via natural language question answering.
-Training and evaluation of multimodal models for the PQA task necessitates comprehensive datasets with scientific textual annotations linked to corresponding protein sequences.
-Currently two PQA-relevant datasets are available.
-While these datasets are suitable for their intended tasks, we found them unfit for the specific purpose of scientific PQA for reasons described below:
+Therefore, the training and evaluation of multimodal models for the PQA task necessitates comprehensive datasets with scientific textual annotations linked to corresponding protein sequences.
+We deemed the following three criteria as essential for a specialized PQA dataset:
   1. 1.
-    ProtDescribe dataset: This dataset (Xu et al., 2023) contains textual annotations from three fields of uncurated SwissProt entries, risking bias and data leakage due to high sequence similarity and over-representation of certain protein families.
-Its use of fields that are not directly inferable from protein sequence (e.g., cellular interaction) limits its utility for PQA.
-Conversely, it misses out on features like catalytic reactions which are often subject of scientific queries.
-Crucially, lacking Q/As, it fails to directly support PQA assessments.
-Therefore, while valuable for enhancing protein representations and text-to-protein generation, as intended by the authors, this dataset exhibits significant limitations for the scientific PQA task.
+    Offers an unbiased representation of known protein sequences, mitigating frequency biases prevalent in existing databases.
   2. 2.
-    PDB-QA dataset: (Guo et al., 2023) With 30 predefined Q/As about 3D structures from the PDB database, while suitable for knowledge retrieval tasks, PDB-QA’s focus on entry-specific details (e.g., submission date, analysis software) renders it ineffective in the context of scientific PQA.
-3.2 Dataset Creation
-To address the gaps discussed in the previous subsection, we created Pika-DS, the first specialized and unbiased PQA dataset, accompanied with respective biologically relevant benchmarks for model training and evaluation.
-We designed this dataset to fulfill three key criteria essential for the training and benchmarking of PQA models:
-  • Offers an unbiased representation of known protein sequences, mitigating frequency biases prevalent in existing databases.
     The dataset ensures that the information associated with each protein sequence is expertly curated, allowing for inference solely based on the protein sequence.
-  • It includes pertinent benchmarking metrics to assess model performance especially in zero-shot settings.
-# total # unique # words # tokens
-Proteins 257 K 252 K - 105.3 M
-Info fields 1.17 M 319 K 10.8 M 29.5 M
-Summaries 1.33 M 861 K 22.5 M 36.4 M
-Questions 1.71 M 98 K 15.7 M 14.9 M
-Answers 1.71 M 783 K 18.4 M 31.8 M
-Table 1: Pika-DS size and token counts. Total and Unique values for Summaries represent the number of sentences.
-Following these criteria, Pika-DS was prepared in three phases:
-  1. 1.
-    We utilized the SwissProt database as a foundational resource.
-SwissProt, the reviewed section of UniProt, features approximately 570,000 detailed annotations of protein sequences spanning all domains of life and viruses.
-For each entry, we extracted an expert-curated list of scientific information covering a wide range of subjects including evolutionary, biochemical and functional properties (Section A.1).
-Removing proteins shorter than 30 amino acids, we obtained 3.7 million information fields for over 548,000 protein sequences.
-  2. 2.
-    UniProt also offers similarity-based clustering of its entries at various sequence similarity thresholds of 100%, 90%, and 50%, known as UniRef clusters.
-For instance, protein sequences that share a minimum of 50% pairwise sequence similarity will be assigned the same UniRef50 cluster.
-Recognizing a strong bias towards more commonly studied protein families in SwissProt (Figure B.1), where nearly 50% of the SwissProt sequences belong to the largest 6% of clusters, we limited Pika-DS to a maximum of two most informative sequences per UniRef50 cluster using a custom algorithm (Section A.2).
-This resulted in a debiased set of over 257,000 protein sequences and 1.17 million information fields.
   3. 3.
-    Lastly, we instructed GPT3.5 to process the information fields for each protein entry to create the Pika-DS’s three main components (example: Table B.1):
+    Supports relevant benchmarking to assess model performance especially in zero-shot settings.
+Since, as discussed in section 2, none of the existing datasets meet all these criteria we created Pika-DS (See sections A.1, A.2, A.3 for details), the first specialized and debiased PQA dataset, accompanied with respective biologically relevant benchmarks for model training and evaluation (summary statistics in Table B.1 and example in Table B.2).
+Briefly, we gathered all SwissProt entries from UniProt database (Consortium, 2022) and extracted an expert-curated list of scientific information fields covering a wide range of relevant properties.
+We then debiased the sequences using a strict 50% similarity threshold.
+Finally, we employed GPT3.5 API to process the information fields for each protein entry using systematically optimized prompts (Sections A.4 and A.5) to create the Pika-DS’s three main components:
   • Summary: A summary of each protein’s functional and biochemical properties, based solely on the provided information excluding the protein’s name.
-    • QAs: Several diverse Q/A pairs for each information field, formatted for LLM training.
-    • Metrics: Single-word answers to a set of predefined scientific questions serving as the ground-truth for our Biochem-Lite benchmarks (Section 3.4).
-    The prompts for all instructions were fine-tuned to ensure GPT3.5’s adherence to using only the given data, summarizing complex biological reactions accurately, and avoiding speculative or ambiguous language.
-Examples in Section A.3.
+  • QAs: Several diverse QA pairs for each information field, formatted for LLM training.
+  • Metrics: Single-word answers to a set of predefined scientific questions serving as the ground-truth for our Biochem-Lite benchmarks (Section 3.2.2).
 Figure 2: Characteristics of PQA dataset.
@@
 Long words are abbreviated (do.=does, mol.=molecule, fun.=function, we.=weight, bel.=belong).
-The final Pika-DS comprises 257,167 protein sequences, carefully selected from 185,128 UniRef50 clusters.
-This dataset is enriched with detailed descriptors for each sequence: a summary statement divided into sentences, multiple scientific Q/A pairs, and the ground-truth answers for our predefined benchmarking questions.
-This dataset encapsulates over 105 million protein sequence tokens, approximately 36.4 million tokens in textual summaries, and nearly 47 million tokens in Q/A pairs (Table 1).
-3.3 Dataset Analysis
-In this subsection, we analyze various facets of our finalized PQA dataset to validate its quality and suitability for training and evaluating multimodal models for the purpose of open-ended scientific enquiries of protein sequences.
-Protein over-representation Analysis:
-We used pre-computed protein embeddings from UniProt database and visualized them using UMAP (Uniform Manifold Approximation and Projection for Dimension Reduction(McInnes et al., 2018)), highlighting the entries that belong to the top 100 largest clusters.
-Visual comparison of distribution of protein embeddings before and after our Uniref50-based filtering confirms a strong reduction of over-represented protein families in Pika-DS while maintaining the rich diversity of the dataset (Figure B.2).
+3.1.1 Pika-DS Quality Control
+The final Pika-DS comprises 257,167 protein sequences, selected from 185,128 UniRef50 clusters.
+This dataset is enriched with detailed descriptors for each sequence: a summary statement divided into sentences, multiple scientific QA pairs, and the ground-truth answers for our predefined benchmarking questions.
+This dataset encapsulates over 105 million protein sequence tokens, approximately 36.4 million tokens in textual summaries, and nearly 47 million tokens in QA pairs (Table B.1).
+To ensure the high quality of Pika-DS and suitability of generated textual annotations for training multimodal LLMs we assessed various aspects of our dataset.
+Human evaluation of GPT3.5 generated annotations:
+Given the potential limitations of GPT3.5 generated annotations (e.g., hallucination), we conducted a thorough human evaluation of the Pika-DS.
+For this, 100 randomly selected examples from Pika-DS were manually evaluated by an expert biochemist.
+Our analysis revealed that out of a total of 1204 summary sentences and QA pairs, 5 were incorrect (0.4%), 21 were of poor quality (1.7%), and 43 were irrelevant (3.6%), resulting in over 94% of the annotations being correct, relevant, and of high quality.
+Furthermore, only 5 out of 660 metric ground truths were found to be incorrect, yielding a 99.2% accuracy in ground truth metric values.
+Furthermore, consultations with two expert biologists confirmed that the QA pairs generated by GPT-3.5 in Pika-DS are relevant and scientifically sound, closely matching what they would derive from the provided input fields.
+They also unanimously agreed that GPT-3.5 questions offer greater diversity compared to their potential queries.
+These results underscore the effectiveness of our prompt optimization approach for generating high-quality labels using GPT3.5 based on SwissProt information fields and is consistent with a recent study that has found LLM-generated captions can exhibit higher diversity than those created by humans, leading to enhanced training of multimodal LLMs (Sharifzadeh et al., 2024).
 Token Count Analysis:
-In creating the Q/A pairs using GPT3.5, we instructed the model to cover a broad spectrum of queries and to produce detailed answers suitable for training other LLMs, anticipating elaborate responses rather than single-word answers.
+In creating the QA pairs using GPT3.5, we instructed the model to cover a broad spectrum of queries and to produce detailed answers suitable for training other LLMs, anticipating elaborate rather than single-word answers.
 This was confirmed by our analysis of token counts in the summary sentences, questions, and answers of the final dataset.
@@
 Word Frequency Analysis:
-To ensure the diversity of content in summary statements as well as Q/A pairs, we examined the frequency of words in the initial five positions of each of these categories.
-As anticipated, common words such as “the” and “it”, in summaries, and interrogatives like “What” and “How” in questions, dominated the first three positions.
-Similarly, answers often began with “the” or “this” as well as “yes/no” structures, followed by the word “protein”.
+To ensure the diversity of content in summary statements as well as QA pairs, we examined the frequency of words in the initial five positions of each of these categories.
+As anticipated, common words such as "the" and "it", in summaries, and interrogatives like "What" and "How" in questions, dominated the first three positions.
+Similarly, answers often began with "the" or "this" as well as "yes/no" structures, followed by "protein".
 However, the remaining positions demonstrated a significant lexical diversity.
 This observation confirmed that our dataset does not exhibit a substantial bias, offering the necessary diversity for effective model training (Figure 2b).
-These analyses underscores the high quality and diversity of the Pika-DS, affirming its suitability for further research in this domain.
-Metric ID Question Example MLP task
-label Self-Pika’s response
-mw MALE
+Protein over-representation Analysis:
+We used pre-computed protein embeddings from UniProt database and visualized them using UMAP (McInnes et al., 2018), highlighting the entries that belong to the top 100 largest clusters.
+Visual comparison of distribution of protein embeddings before and after our Uniref50-based filtering confirms a strong reduction of over-represented protein families in Pika-DS while maintaining the rich diversity of the dataset (Figure B.2).
+3.2 Pika Benchmarks
+In this section, we elaborate on our biochemically-focused benchmarking methodologies tailored for evaluating the scientific accuracy of multimodal PQA models.
+3.2.1 Motivation & Design Criteria
+Conventional linguistic metrics like BLEU and ROUGE, while useful in general linguistic contexts, often fall short in assessing scientific correctness and show poor correlation with human judgment (Mathur et al., 2020).
+As a result these metrics are inadequate for assessing the performance of multimododal PQA models.
+Therefore, going beyond standard linguistic evaluations, we designed a purpose-built benchmarking approach incorporating a set of predefined, biochemically-significant questions, that are specifically selected to:
+  1. 1.
+    Reflect the biochemical properties of proteins, ensuring that a model’s accuracy in these responses is indicative of its effectiveness in broader scientific enquiries.
+  2. 2.
+    Span a spectrum of complexity, from straightforward information extraction with minimal linguistic intricacy to advanced linguistic reasoning for identifying pertinent information.
+In consultation with domain experts, we identified five core protein properties at distinct difficulty levels that form the basis for our scientific benchmarking questions: molecular weight (mw), co-factor binding, sub-cellular localization, protein domains, and enzymatic reaction.
+Alongside these benchmarking questions, we also require a robust strategy to assess the correctness of open-ended answers to them.
+Four possible assessment strategies could be imagined: (1) Exact matching of statements, (2) Keyword comparison, (3) Comparison using stronger LLMs (xLLMs), and (4) Human evaluation.
+While exact statement matching is useful in some research domains, it lacks the rigour required for assessing scientific accuracy.
+On the other hand, human evaluation for specialized domains is extremely costly at large scales.
+However, we can use human supervision to ensure the quality of comparisons conducted by xLLMs.
+As a result we selected keyword comparison and use of xLLMs guided by human experts for Pika framework.
+3.2.2 Two-Tiered Benchmarking System
+The use of xLLMs for response evaluation is more accurate but computationally prohibitive as compared to keyword comparison, which pragmatically balances scientific accuracy with computational efficiency.
+This balance of performance vs efficiency motivated us to design a two-tiered benchmarking system comprised of a light-weight benchmark employing keyword comparison for efficiency, and a rigorous xLLM-based benchmark for scientific fidelity.
+Metric ID Question Example Label Example Response
+mw MALE $\downarrow$
 What is the molecular weight of
@@
 54988 KDa.
-log value
-exact cofactor What is a cofactor of this protein? Zn(2+) The cofactor of this protein is Zn(2+). -
-is_enzyme F1
+exact cofactor $\uparrow$ What is a cofactor of this protein? Zn(2+) The cofactor of this protein is Zn(2+).
+is_enzyme F1 $\uparrow$
 Can this sequence be considered
 an enzyme?
-True Yes classification
-location F1 Where is this protein located? Membrane
+True Yes
+location F1 $\uparrow$ Where is this protein located? Membrane
 The sub-cellular location of this protein
 is the cell membrane.
-classification
 binary Q
-mean F1
+mean F1 $\uparrow$
 Is this a membrane protein?
@@
 No
-Table 2: Questions and example responses for Biochem-Lite (Section A.4 provides further details on motivation and score computation for each metric).
-Self-Pika is our best performing model.
-Underlined values indicate the extracted entities for score calculation.
-3.4 Benchmarks
-In this section, we elaborate on our biochemically-focused benchmarking methodologies tailored for evaluating the scientific accuracy of multimodal PQA models.
-Motivation & Design Criteria:
-Conventional linguistic metrics like BLEU and ROUGE, while useful in general linguistic contexts, often fall short in assessing scientific correctness and show poor correlation with human judgment (Mathur et al., 2020).
-As a result these metrics are inadequate for assessing the performance of multimododal PQA models.
-Therefore, going beyond standard linguistic evaluations, we designed a purpose-built benchmarking approach incorporating a set of predefined, biochemically-significant questions, that are specifically selected to:
-  1. 1.
-    Reflect the biochemical properties of proteins, ensuring that a model’s accuracy in these responses is indicative of its effectiveness in broader scientific enquiries.
-  2. 2.
-    Span a spectrum of complexity, from straightforward information extraction with minimal linguistic intricacy to advanced linguistic reasoning for identifying pertinent information.
-Two-Tiered Benchmarking System:
-Alongside the benchmarking questions for the PQA task, we also require a robust strategy to assess the correctness of open-ended answers to them.
-Three possible assessment strategies could be imagined: (1) Exact matching of statements, (2) Relevant entity extraction, and (3) Comparison using stronger LLMs (xLLMs).
-While exact statement matching is useful in some research domains, it lacks the rigour required for assessing scientific accuracy, as such we selected entity extraction and xLLM comparison.
-While xLLM comparison, although ideal, is computationally prohibitive, entity extraction pragmatically balances scientific accuracy with computational efficiency.
-This balance of performance vs efficiency motivated us to design a two-tiered benchmarking system comprised of a light-weight benchmark employing entity extraction for efficiency, and a rigorous xLLM-based benchmark for scientific fidelity.
-Metric ID Question Ground truth Self-Pika’s response
+Table 1: Questions and example responses for Biochem-Lite (See A.6 for our motivation for each metric).
+Underlined indicates extracted entities for score calculation.
+$\uparrow$: Higher values better.
+$\downarrow$: Lower values better.
+Metric ID Question Ground Truth Example Response
 Reaction
@@
 pyridoxal 5’-phosphate and magnesium ions.
-Table 3: Questions and example responses for Biochem-ReAct. Self-Pika is our best performing model.
+Table 2: Questions and example responses for Biochem-ReAct.
+Presented in order of difficulty, with Reaction being the most difficult question for a multimodal LLM to answer based on a protein sequence.
   1. 1.
-    Biochem-Lite: Pika’s light-weight benchmarking involves a set of pre-defined, scientifically relevant questions with simple answers extracted for each protein using GPT3.5 during Pika-DS creation.
+    Biochem-Lite: Pika’s light-weight benchmarking involves a set of pre-defined, scientifically relevant questions with simple answers extracted for each protein using GPT3.5 during Pika-DS creation (Table 1).
 These questions are designed to cover a range of biochemical and functional properties (e.g.
-binding to cofactros or cellular localization) while evaluating multimodal PQA performance at various levels of (1) information extraction from protein embeddings, (2) cross-modal information processing by the LLM, or (3) the LLM’s ability to generate relevant responses (Table 2 and Section A.4).
-Although the ground truth answers to these questions are single-worded, for evaluation, questions are presented in free-form, and the model provides an open-ended response.
-This response is then processed for rule-based entity extraction and scoring using adequate metrics.
-    • For mw MALE, the mean-absolute log error (MALE) of the predicted molecular weight (MW) is:
-      $MALE=\frac{1}{N}\sum_{i=1}^{N}|\log_{10}(\hat{MW}_{i})/(MW_{i})|$ (1)
-      where $\hat{MW}_{i}$ is the predicted MW, $MW_{i}$ is the ground truth MW, and $N$ is number of examples.
-    • For exact cofactor, the score is computed as:
-      $\text{Score}_{\text{exact}}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{1}(\exists w\in R_{i}:w\in GT_{i})$ (2)
-      where $R_{i}$ is the set of words in response $i$, $GT_{i}$ is the set of ground truth cofactor words for protein $i$, and $\mathbb{1}$ is the indicator function that is 1 for exactly one match and 0 otherwise.
-    • For location F1, the F1 score is defined as:
-      $F1=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (3)
-      The predicted class assignments for the calculation of Precision and Recall are based on the exclusive presence of correct labels (membrane, nucleus, or mitochondrion) in the generated response.
-Responses lacking or containing multiple labels receive the none class label.
-    • For is_enzyme F1, the F1 score is used as defined in Equation 3.
-Responses are classified as True or False based on the exclusive presence of yes or no, respectively.
-Any deviation is labeled as none.
-    • For binary Q mean F1, we have:
-      $F1_{\text{binary Q mean}}=\frac{1}{3}\sum_{q=1}^{3}F1_{q}$ (4)
-      where $F1_{q}$ is the F1 score for each binary question with class assignments similar to the is_enzyme F1.
+binding to cofactros or cellular localization) while evaluating multimodal PQA performance at various levels (i.e., information extraction from protein embeddings, cross-modal information processing by the LLM, or the LLM’s ability to generate relevant responses).
+Although the ground truth answers to these questions are single-worded, for evaluation, the questions are presented in free-form, and the model provides open-ended responses, which is processed via rule-based entity extraction and is scored using adequate metrics as detailed in Section A.7.
   2. 2.
     Biochem-ReAct: While light-weight benchmarking questions offer a general overview of model’s performance in various scientific aspects, they do not represent real-world scenarios where the scientific fidelity of open-ended responses is vital.
-Therefore, we also selected three biochemical questions surrounding reaction & activity of the proteins that can be subjects of scientific enquiry into novel protein sequences (Table 3).
-The complexity of the responses to these questions necessitates assessment using advanced LLMs such as GPT3.5 and GPT4, therefore imposing a significant computational cost.
-As a result, Biochem-ReAct is mainly suitable for end-point benchmarking of the best-performing models.
+Therefore, we also selected three biochemical questions at three distinct difficulty levels surrounding Reaction & Activity of the proteins that can be subjects of scientific enquiry into novel protein sequences (Table 2).
+The complexity of the responses to these questions necessitates assessment using advanced LLMs such as GPT4.
+To ensure the high reliability of these assessments we performed an iterative prompt optimization with human feedback following the approach described in Section A.4.
 Overall, our two-tiered benchmarking system establishes a balanced and robust framework for evaluating multimodal PQA models, ensuring both scientific rigor and computational feasibility.
-4 Baselines & Methodology
-In this section we outline the baselines, architecture and the training strategy for our proposed solutions to the task of scientific PQA.
-All baseline models and architectures were implemented using PyTorch Lightning, and the complete codebase is accessible on github.com/EMCarrami/Pika.
-4.1 Baselines
-The prior research in the subject of multimodal text-protein models has primarily been focused on improving protein embeddings with textual data or for text-guided protein sequence generation (Xu et al., 2023; Zhang et al., 2022).
-The use of PQA in other studies (Guo et al., 2023) has been limited to knowledge retrieval, which do not provide comparable benchmarking opportunities for scientific PQA, mainly due to the high presence of non-scientific content in the PDB-QA dataset (e.g., author names).
-Therefore, in the absence of relevant state-of-the-art, here we define four baselines specifically for scientific PQA in the context of Biochem-Lite benchmarking questions (Section 3.1).
-  • Random Baseline: Selects a random answer for each question from the pool of all possible answers to the same question within the evaluated dataset/split.
-  • MLP: An MLP on protein sequence embeddings of the PLM (Section A.5).
-Applies to classification or value prediction tasks in Biochem-Lite (Table 2).
-It sets an upper limit for the information content of PLM embeddings, with any improvements potentially indicating generalization due to the knowledge encapsulated in the LLM.
-  • LLM Only: Pre-trained LLM’s base response to Biochem-Lite questions. Phi-2 was used.
-  • Pika w/o PLM: Mirrors Self-Pika’s architecture and training, substituting the PLM with simple token embeddings of the protein sequences.
-It defines the base performance for the Biochem-Lite benchmarks (except for “mw MALE”).
+3.3 Pika Baselines
+As discussed in Section 2, previous PQA studies have either been limited to knowledge retrieval or have ignored the evolutionary context of protein sequences.
+Combined with the use of unsuitable metrics such as BLEU, these studies do not provide relevant benchmarking opportunities for scientific PQA.
+Furthermore, unlike natural multimodal LLMs (e.g., Vision Question Answering) where human-baselines are possible, in PQA it is not possible to determine the upper-bound of performance by comparison to expert evaluations, because humans are unable to extract any information from protein sequences.
+Therefore, in the absence of relevant baselines or human evaluation for PQA, we introduce two multimodal PQA-LLMs and define various lower- and upper-bound baselines.
 Figure 3: Schematic representation of Cross- and Self-Pika architectures for the scientific PQA task.
@@
 Only the adapter and cross-attention modules (both in green) are trained.
-Models Biochem-Lite Biochem-ReAct
+3.3.1 Pika Models
+We propose two robust multimodal architectures for PQA tasks, drawing inspiration from successful Vision Language Model (VLM) strategies, replacing the vision-encoder with a protein language model (PLM) (Fig.
+3).
+  1. 1.
+    Cross-Pika is inspired from Flamingo (Tsimpoukelli et al., 2021) and Prismer (Liu et al., 2023).
+It uses multiple independent learnable adapters, each creating distinct protein latent embeddings for each transformer layer of the LLM.
+These embeddings are injected into the LLM using a gated cross-attention mechanism before each native self-attention layer.
+At the cost of increased complexity, this design allows for a nuanced and layer-specific modulation of the LLM, potentially enabling it to process complex protein-related information more effectively, which could aid the model in answering biochemically intricate PQA queries.
+  2. 2.
+    Self-Pika is based on the architecture proposed in Frozen (Tsimpoukelli et al., 2021) and uses a single learnable adapter to transform protein sequence embeddings into latent embeddings compatible with language token embeddings.
+The transformed embeddings, concatenated at the beginning of the LLM’s initial token embeddings, allow the protein latent embeddings to influence the LLM’s response through its internal self-attention mechanisms.
+This approach effectively creates soft-prompts conditioned on the input protein sequence.
+This architecture simplifies the integration process, reducing computational overhead while allowing for the incorporation of essential protein characteristics into the LLM’s response.
+Considering the diversity in protein lengths, we opted for the Perceiver architecture (Jaegle et al., 2021) as the learnable adapter for both models.
+This choice standardizes the transformation of protein embeddings into a consistent number of latent embeddings for seamless cross-modality information transfer.
+In both architectures we keep the pre-trained LLM and PLM frozen, only incorporating trainable "adapter" modules that extract relevant latent representations from PLM’s output (protein sequence embeddings) and integrate them into the LLM’s transformer architecture, facilitating seamless transfer of information across modalities.
+In this work we use ESM2 (Lin et al., 2022) as the pre-trained PLM and Phi-2 (Microsoft, 2023) as the pre-trained LLM (In some experiments GPT2 (Radford et al., 2019) was used, where indicated).
+3.3.2 Lower-bound Baselines
+  • Random Baseline: Selects a random answer for each question from the pool of relevant answers.
+  • LLM Only: Pre-trained Phi-2 model to determine LLM’s base response to Biochem-Lite questions.
+  • Pika w/o PLM: Mirrors Self-Pika’s architecture and training, substituting the PLM with simple token embeddings of the protein sequences.
+This effectively reduces the information to amino acid content only, ignoring the context and order of the sequence.
+3.3.3 Upper-bound Baselines
+  • MLP: An MLP on protein sequence embeddings of ESM2 (Details in Section A.8).
+It defines the upper limit for the information content of ESM2 embeddings, with any performance beyond this by a model potentially indicating generalization due to the knowledge encapsulated in the LLM.
+  • BLAST: Considering the evolutionary context of protein sequences, for any queried sequence, we identify the closest related protein sequence in the training data using BLAST (Camacho et al., 2009) and return the respective information as the prediction.
+This is a known strong upper-bound baseline, as homology is a strong predictor of protein’s properties.
+4 Experiments & Results
+In this section we share our key experimental results. Training details can be found in Sections A.9 and A.10.
+Figure 4: Evaluating the effectiveness of Biochem-Lite vs traditional linguistic metrics for scientific accuracy of PQA.
+Statistical significance is determined through a one-tailed paired t-test across three randomly seeded data subsets and model training (significance guide: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$).
+4.1 Validation of Biochem-Lite
+First, we assessed the reliability of Biochem-Lite metrics in assessing PQA models performance.
+To We compared a simple Pika model (Self-Pika with ESM2 + GPT2) against its truncated version, Pika w/o PLM (Section 3.3.2), where ESM2 model was replaced with simple protein token embeddings.
+Considering that the removal of ESM2 should eliminate all functional information, leaving behind only the amino acid content, we expect that the truncated model must perform significantly worse than the original model.
+Therefore, metrics that fail to show a significant difference between these two would be unsuitable for assessing PQA models.
+Our results indicate that Biochem-Lite metrics are significantly better in Self-Pika vs Pika w/o PLM, while all ROUGE metrics fail to highlight any differences (Fig.
+4).
+The behaviour of the mw MALE metric is expected as the size of a protein only depends on its amino acid content.
+This finding confirms the inadequacy of traditional linguistic metrics for scientific PQA task, demonstrating the utility and importance of our Biochem-Lite metrics.
+4.2 Zero-shot PQA:
+We define zero-shot PQA as answering questions about unseen protein sequences.
+Considering the debiased nature of Pika-DS, this can easily be achieved by splitting the data based on UniRef50 ID of each sequence, thereby ensuring that the validation/test sequences have no more than 50% sequence similarity with any of the training sequences.
+While this ensures the unseen sequence criteria, when considering the evolutionary context of proteins, the 50% similarity threshold is a lenient cutoff.
+This is because homologous proteins may exhibit as low as 15% sequence similarity (Leander et al., 2022).
+As a result, on UniRef50-based splits, model performance will be a combination of generalization based on the sequence and model’s ability to identify distant evolutionary relations to training data.
+While such behaviour may be desired for some PQA applications, to extend the scope, we created a more stringent splitting strategy that focuses the evaluation of model’s generalization primarily on the basis of proteins sequences without the impact of evolutionary relationships.
+To achieve this we minimized the the evolutionary connection of validation/test sequences to the training data.
+This was achieved by first grouping sequences into evolutionary groups (EvoGroups) (See Section A.11 for details) followed by splitting the data based on their EvoGrop assignments, ensuring that sequences in validation and test do not have closely related evolutionary counterparts in the training data.
+Naturally, we expect that the BLAST baseline should show a strong performance on UniRef50-based splits, while its performance should diminish on EvoGroup-based splits.
+Baseline or Model Biochem-Lite Biochem-ReAct
 mw
@@
 mF1
-pplx Reaction Domains Cofactor
-Random 0.35 0.13 0.49 0.34 0.46 - - - -
-Baselines MLP 0.07 - 0.86 0.89 - - - - -
-Phi-2 only 4.53 0.03 0.34 0.00 0.31 - - - -
-Pika w/o PLM 0.02 0.21 0.7 0.47 0.21 2.30 0.01 0.32 0.82
-Cross-Pika (Q) 0.02 0.47 0.89 0.77 0.13 1.69 0.36 0.60 0.94
-Our Models Self-Pika (Q) 0.04 0.54 0.89 0.76 0.44 1.98 0.58 0.79 0.97
-Self-Pika (Q+S) 0.04 0.57 0.88 0.76 0.57 2.23 0.53 0.75 0.96
-Table 4: Performance of Pika models on Biochem-Lite and Biochem-ReAct metrics.
-Bold values indicate the best score.
-Standard deviation for all scores $<0.05$ except for those indicated by underline.
+Reaction Domains Cofactor
+Random 0.35 (0.01) 0.13 (0.02) 0.49 (0.00) 0.34 (0.02) - - - -
+Lower-bound Phi-2 only 4.53 (0.01) 0.03 (0.01) 0.34 (0.00) 0.00 (0.00) 0.31 (0.00) - - -
+Pika w/o PLM 0.02 (0.01) 0.21 (0.01) 0.7 (0.01) 0.47 (0.01) 0.21 (0.08) 0.01 0.32 0.82
+Upper-bound MLP 0.07 (0.01) - 0.86 (0.01) 0.89 (0.03) - - - -
+BLAST - 0.71 (NA) 0.88 (NA) 0.93 (NA) - 0.94 0.89 0.99
+Pika Models Cross-Pika 0.02 (0.00) 0.47 (0.15) 0.89 (0.02) 0.77 (0.01) 0.13 (0.02) 0.36 0.60 0.94
+Self-Pika 0.04 (0.02) 0.54 (0.04) 0.89 (0.01) 0.76 (0.02) 0.44 (0.04) 0.58 0.79 0.97
+Table 3: Performance of Pika models on Uniref50 splits.
+Bold values indicate the best score, underline indicates second best score.
+Values in ( ) indicate standard deviation.
 Biochem-Lite results on val set and from three different seeded training.
-Biochem-ReAct results on test set and a single seed training.
-“binary mF1” = “binary Q mean F1”, pplx = perplexity, Q = training with QA data only, Q+S = training with QA and Summary data.
-4.2 Models
-We propose two robust multimodal architectures for PQA tasks, drawing inspiration from successful VLM strategies like Frozen (Tsimpoukelli et al., 2021), Flamingo (Tsimpoukelli et al., 2021) and Prismer (Liu et al., 2023).
-In both architectures we keep the pre-trained LLM and PLM frozen, only incorporating trainable “adapter” modules that extract relevant latent representations from PLM’s output (protein sequence embeddings) and integrate them into the LLM’s transformer architecture, facilitating seamless transfer of information across modalities.
-  • Cross-Pika uses multiple independent learnable adapters, each creating distinct protein latent embeddings for each transformer layer of the LLM.
-These embeddings are injected into the LLM using a gated cross-attention mechanism before each native self-attention layer (Fig.
-3).
-At the cost of increased complexity, this design allows for a nuanced and layer-specific modulation of the LLM, potentially enabling it to process complex protein-related information more effectively, which could aid the model in answering biochemically intricate PQA queries.
-  • Self-Pika uses a single learnable adapter to transform protein sequence embeddings into latent embeddings compatible with language token embeddings.
-The transformed embeddings, concatenated at the beginning of the LLM’s initial token embeddings, allow the protein latent embeddings to influence the LLM’s response through its internal self-attention mechanisms.
-This approach effectively creates soft-prompts conditioned on the input protein sequence (Fig.
-3).
-This architecture simplifies the integration process, reducing computational overhead while allowing for the incorporation of essential protein characteristics into the LLM’s response.
-Considering the diversity in protein lengths, we opted for the Perceiver architecture (Jaegle et al., 2021) as the learnable adapter for both models.
-This choice standardizes the transformation of protein embeddings into a consistent number of latent embeddings for seamless cross-modality information transfer.
-Optionally, predefined placeholder tokens can also be added to each question to signaling the LLM about the presence of the protein sequence.
-In this work we used ESM2 (Lin et al., 2022) as the pre-trained PLM and Phi-2 (Microsoft, 2023) or GPT2 (Radford et al., 2019) as the pre-trained LLM.
-Our best performing model is based on the Self-Pika architecture, with ESM2-L (esm2_33t_650M) and Phi-2 pre-trained models.
-4.3 Training
-We devised a simple training strategy for PQA models, keeping both the PLM and LLM frozen and optimizing for the causal language model loss with AdamW.
-Hyperparameters were optimised following a greedy search as detaile in section A.6.
-All training was performed on a single A100-80GB or H100-80GB Nvidia GPU.
-Unless otherwise stated, the QA section of the Pika-DS was used for the training of all Pika models.
-Examples were split to train, validation and test set based on the UniRef50 cluster of their respective protein sequences.
-As a result, protein sequence from the evaluation sets (validation and test) are never present in training data.
+Biochem-ReAct results on test set and a single seed.
+4.2.1 Performance on UniRef50 Splits
+Table 3 summarizes the benchmarking results of Pika models on a random UniRef50-based split in comparison with various baselines.
+It is evident that both Pika architectures outperform all lower-bound baselines, indicating successful cross-modality information transfer to the LLM.
+Notably, our top model surpasses the upper-bound MLP baseline on key Biochem-Lite benchmarking questions, underscoring the potential for generalization via the LLM’s knowledge.
+However, as expected, the strong BLAST baseline outperforms all models in most Biochem-Lite and all Biochem-ReAct metrics.
+Nevertheless, these results indicate that our Pika models based on the modestly sized Phi-2 LLM, are able to comprehend distant evolutionary relations form the context of sequence embeddings.
+Baseline or Model Biochem-Lite Biochem-ReAct
+exact
+cofactor
+is_enzyme
+F1
+location
+F1
+Reaction Domains Cofactor
+BLAST Baseline 0.21 0.72 0.51 0.52 0.49 0.94
+Self-Pika Model 0.57 0.85 0.69 0.09 0.54 0.87
+Table 4: Performance of Pika model on EvoGroup splits. Bold values indicate the best score. Biochem-Lite results on val set and Biochem-ReAct results on test set both with a single seed.
+4.2.2 Performance on EvoGroup Splits
+To assess generalization by Pika models, in isolation from the effects of evolutionary relationship of proteins, we performed Self-Pika training on EvoGroup splits and compared the results with BLAST baseline on these splits (Table 4).
+As expected, our results show a significant reduction in performance of BLAST baseline, specially in more complex metrics such as Domains and Reaction, and to a lower extent in simpler metrics such as Cofactor.
+This is because, cofactor binding, for instance, relies on small protein motifs that are more conserved across distant evolutionary relations.
+Crucially, Self-Pika, except on the most complex Reaction metric, retained the majority of its performance on EvoGroup splits, allowing it to surpass the performance of BLAST baseline in most metrics.
+These observations indicate that multimodal PQA models are capable of inferring functional properties both based on the sequence of proteins as well as using distant evolutionary relations.
+4.2.3 Correlation of Biochem Metrics
+The Self-Pika model demonstrated strong performance across all three Biochem-ReAct questions on the UniRef split, exceeding the without PLM baseline by 57%, 47%, and 15% for correctly identifying Reactions, Domains, and Cofactors, respectively, for previously unseen protein sequences.
+This prompted us to study the correlation of Biochem-Lite metrics to Biochem-ReAct metrics.
+Comparing Biochem-Lite results against Biochem-ReAct scores for 12 high-performing checkpoints with varied configurations revealed that while traditional linguistic metrics, including perplexity, fall short in predicting the real-world efficacy of PQA models, two Biochem-Lite questions, "exact cofactor recall" and "binary Q mean F1", emerged as strong indicators of multimodal PQA models’ real-world performance based on Biochem-ReAct metrics (Fig.
+B.3).
+4.2.4 PQA Learning without Questions
+To understand the significance of QA pairs in the training set, we conducted training under two stringent conditions.
+The first involved using only summary sentences as labels for proteins without any QAs.
+The second condition included summary sentences and a single Control Question, "Is this a real protein?".
+During training, we randomly shuffled the tokens within the sequence of half the proteins, setting the expected answer to this question as No (for summary labels all protein sequences remained unchanged).
+This aims to train the model to understand the task of question answering, with summaries providing the scientific context for learning.
+Remarkably, introducing the Control Question, even when no other questions were present during training, significantly enhanced performance (Table 5).
+Additionally, we observed that the performance of the model on the Control Question follows a similar pattern as other metrics (Table B.3).
+These observation suggest that the current bottleneck in PQA performance is likely the LLM and extending this work to larger LLMs could further improve performance and generalization capabilities.
 Training Mode Biochem-Lite Biochem-ReAct
@@
 self-Pika (Q+S) 0.04 0.57 0.57 0.53 0.75 0.96
-Table 5: Perfomance of Self-Pika in the absence of Q/As during training.
+Table 5: Perfomance of Self-Pika in the absence of QAs during training.
 Bold values indicate best score.
 Standard deviation for all scores $<0.05$ except for those indicated by underline.
-S = Summary, Ctrl Q = Control Question, QA = Q/A pairs.
-5 Experiments & Results
-We start this section by validating our benchmarking strategy, followed by the zero-shot evaluation of best models.
-In the end, we present the results of our ablation study, highlighting the key factors in scientific PQA model performance.
-5.1 Benchmarking Results
-Biochem-Lite Results:
-Table 4 summarizes the benchmarking results of Pika models in comparison with various baselines.
-It is evident that both Self- and Cross-Pika architectures outperform all lower-bound baselines, indicating successful cross-modality information transfer to the LLM.
-Notably, our top models also surpass the upper-bound MLP baseline on key Biochem-Lite benchmarking questions, underscoring the potential for generalization through the LLM’s knowledge.
-Additionally, we assessed the reliability of ROUGE scores for evaluating PQA model performance.
-ROUGE scores remained consistent, regardless of whether Self-Pika utilized simple protein token embeddings (w/o PLM) or embedded protein sequences (Figure B.3).
-This finding confirms the inadequacy of traditional linguistic metrics for the scientific PQA task, thus demonstrating the utility of our Biochem-Lite metrics.
-Biochem-ReAct Results:
-Our Biochem-ReAct benchmarking, aimed at measuring scientific fidelity, yielded promising results (Table 4).
-The Self-Pika model demonstrated strong performance across all three Biochem-ReAct questions, exceeding the without PLM baseline by 57%, 31%, and 15% for correctly identifying Reactions, Domains, and Cofactors, respectively, for previously unseen protein sequences.
-Additionally, comparing Biochem-Lite results against Biochem-ReAct scores for 12 high-performing checkpoints with varied configurations revealed that while traditional linguistic metrics, including perplexity, fall short in predicting the real-world efficacy of PQA models, two Biochem-Lite questions—“exact cofactor recall” and “binary Q mean F1”—emerged as strong indicators of multimodal PQA models’ real-world performance (Figure B.4).
-Overall, these results validate our benchmarking strategy and establish Self-Pika as the initial state-of-the-art for the task of scientific PQA.
-5.2 Zero-shot PQA:
-We define zero-shot PQA as answering novel questions about unseen protein sequences.
-Given the unbiased nature of Pika-DS and data splits based on UniRef50 IDs, our validation and test sets largely comprise unseen protein sequences.
-Moreover, the “binary Q” component in Biochem-Lite consists of questions never included in the Pika-DS, suggesting our benchmarking broadly represents zero-shot PQA tasks.
-This section delves into stricter zero-shot scenarios to better understand PQA model behavior.
-PQA Learning without Questions:
-To understand the significance of Q/A pairs in the training set, we conducted training under two stringent conditions.
-The first involved using only summary sentences as labels for proteins without any Q/As.
-The second condition included summary sentences and a single Control Question, “Is this a real protein?”.
-During training, we randomly shuffled the tokens within the sequence of half the proteins, setting the expected answer to this question as No (for summary labels all protein sequences remained unchanged).
-This aims to train the model to understand the task of question answering, with summaries providing the scientific context for learning.
-Remarkably, introducing the Control Question, even when no other questions were present during training, significantly enhanced performance (Table 5).
-Additionally, we observed that the performance of the model on the Control Question follows a similar pattern as other metrics (Table 6).
-These observation suggest that the current bottleneck in PQA performance is likely the LLM and extending this work to larger LLMs could further improve performance and generalization capabilities.
-Random
-LLM
-only
-self-Pika
-w/o PLM
-self-Pika
-F1 Score 0.49 0.00 0.43 0.99
-Accuracy 0.50 0.00 0.51 0.99
-Table 6: Performance on Control Question.
+S = Summary, Ctrl Q = Control Question, QA = QA pairs.
+4.2.5 PQA Learning with Novel Proteins
+In training multimodal Pika models, we ensured that proteins similar to those in the training dataset were excluded from the test and validation sets, effectively creating a zero-shot scenario for our benchmarking.
+Given that the ESM2 model was pre-trained on a vast dataset of proteins, there was a potential concern that its embeddings might be influenced by previously encountered sequences.
+To mitigate this, we utilized the evaluation split reported for ESM2’s pre-training, categorizing our benchmarking results into proteins seen and unseen by ESM2.
+The analysis revealed no significant performance difference between these conditions, affirming that Pika’s zero-shot capabilities are not compromised by the prior knowledge embedded in ESM2.
+This finding underscores the robustness of Pika models in genuine zero-shot PQA tasks, independent of ESM2’s pretraining exposure.
+4.2.6 Ablation Studies
+To identify key contributing factors, we compared model performance across various dimensions.
+Most notably, reducing the size of the LLM from Phi-2, which has 2.8 billion parameters, to GPT-2 Medium, with 355 million parameters, resulted in a 20% decrease in the key "exact cofactor" metric (Table 6).
+However, the size of the ESM2 model did not seem to have a significant impact, with performance remaining largely similar between ESM2-S, ESM2-M, and ESM2-L (8M, 65M and 350M parameters, respectively).
+This observation could be attributed to the richness of representations in the Pre-trained PLMs, potentially highlighting a bottleneck in the LLM or the quality and quantity of the data (Table 6).
+Concordantly, the comparison of results from the zero-shot experiment also confirms that with a lower volume of data, a tailor-made dataset could yield better results.
+However, as the dataset size increases—a scenario analogous to the image domain—this factor becomes less crucial.
+These findings suggest that while LLM capacity significantly influences PQA model performance, the sophistication of sequence embeddings provided by ESMs reaches a point of diminishing returns, underscoring the importance of focusing on LLM enhancements and data quality for future improvements.
 LLM PLM
@@
 Table 6: Effect of LLM and PLM size on Self-Pika models’ performance. Standard deviation for all scores $<0.05$ except for those indicated by underline.
-PQA Learning with Novel Proteins:
-In training multimodal Pika models, we ensured that proteins similar to those in the training dataset were excluded from the test and validation sets, effectively creating a zero-shot scenario for our benchmarking.
-Given that the ESM2 model was pre-trained on a vast dataset of proteins, there was a potential concern that its embeddings might be influenced by previously encountered sequences.
-To mitigate this, we utilized the evaluation split reported for ESM2’s pre-training, categorizing our benchmarking results into proteins seen and unseen by ESM2.
-The analysis revealed no significant performance difference between these conditions, affirming that Pika’s zero-shot capabilities are not compromised by the prior knowledge embedded in ESM2.
-This finding underscores the robustness of Pika models in genuine zero-shot PQA tasks, independent of ESM2’s pretraining exposure.
-5.3 Ablation Studies
-To identify key contributing factors, we compared model performance across various dimensions.
-Most notably, reducing the size of the LLM from Phi-2, which has 2.8 billion parameters, to GPT-2 Medium, with 355 million parameters, resulted in a 20% decrease in the key “exact cofactor” metric (Table 7).
-However, the size of the ESM2 model did not seem to have a significant impact, with performance remaining largely similar between ESM2-S, ESM2-M, and ESM2-L (8M, 65M and 350M parameters, respectively).
-This observation could be attributed to the richness of representations in the Pre-trained PLMs, potentially highlighting a bottleneck in the LLM or the quality and quantity of the data (Table 7).
-Concordantly, the comparison of results from the zero-shot experiment also confirms that with a lower volume of data, a tailor-made dataset could yield better results.
-However, as the dataset size increases—a scenario analogous to the image domain—this factor becomes less crucial.
-These findings suggest that while LLM capacity significantly influences PQA model performance, the sophistication of sequence embeddings provided by ESMs reaches a point of diminishing returns, underscoring the importance of focusing on LLM enhancements and data quality for future improvements.
-6 Conclusion
+5 Conclusion & Impact
 We’ve established a pioneering framework for zero-shot PQA, presenting a significant step forward in the application of LLMs for scientific enquiry.
-The introduction of specialized datasets and biologically relevant benchmarks underpins future explorations at the intersection of computational biology and artificial intelligence.
-Key insights from our zero-shot evaluations and ablation studies highlight the crucial role of LLMs in multimodal PQA performance, underscoring the importance of the knowledge depth and linguistic capabilities of LLMs.
-In essence, this work not only sets a new standard for scientific PQA but also opens the door for the research community to further refine and expand upon our methodologies.
-Looking ahead, leveraging larger, more diverse LLMs may offer substantial gains in model generalization, driving us towards the goal of automating accurate scientific enquiry.
-7 Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning and Biology.
-There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-Acknowledgements
+The introduction of our specialized datasets and biologically relevant benchmarks underpins future explorations at the intersection of computational biology and artificial intelligence.
+Key insights from our zero-shot evaluations and ablation studies highlight the crucial role of LLMs in multimodal PQA performance, while underscoring the importance of considering evolutionary relation in assessing performance.
+Since, for Pika-DS we employed GPT3.5 to create synthetic annotations, given the limitations of LLMs in generating large-scale synthetic datasets, we endeavored to minimize the inclusion of harmful content in Pika-DS through prompt optimization and manual evaluation.
+Nonetheless, due to the dataset’s extensive size, there’s a slight possibility that unintended harmful content might still be present.
+Our Pika-based pre-trained models are derived from the publicly accessible and unmoderated Phi-2 LLM.
+Thus, all cautions, restrictions, and notes associated with Phi-2 (Microsoft, 2023) are applicable to our models.
+Looking ahead, leveraging larger, more diverse LLMs may offer substantial gains in model generalization, driving us towards the goal of automating accurate scientific enquiry into proteins.
+6 Reproducibility & Accessibility
+All code and data used for creation of data, model training and baselines are publicly accessible on www.github.com/EMCarrami/Pika.
+The Pika framework is specifically designed for question answering related to protein sequences.
+With scientists having identified nearly 0.25 billion protein sequences, and functional annotations available for fewer than a million, our framework offers significant potential for research into these largely unexplored proteins.
+While our efforts are directed towards scientific research, we recognize the potential risk of misuse by individuals aiming to create or identify harmful substances.
+We strictly prohibit using our dataset and framework for any illegal activities or actions that could harm individuals or society.
+Acknowledgments and Disclosure of Funding
 We would like to thank Dr Aida Nematzadeh for critical advice and Maryam Karaminejad Ranjbar for graphic design.
+We would also like to sincerely thank Relation Therapeutics for providing us with the necessary compute for training the final version of Pika.
+This work was primarily supported by personal funding and the authors declare no conflict of interest.
 Appendix A Supplementary Methods
-A.1 Data collection from UniProt
+A.1 Creation of Pika-DS
+Pika-DS was prepared in three phases:
+  1. 1.
+    We utilized the SwissProt database as a foundational resource.
+SwissProt, the reviewed section of UniProt, features approximately 570,000 detailed annotations of protein sequences spanning all domains of life and viruses.
+For each entry, we extracted an expert-curated list of scientific information covering a wide range of subjects including evolutionary, biochemical and functional properties (Section A.2).
+Removing proteins shorter than 30 amino acids, we obtained 3.7 million information fields for over 548,000 protein sequences.
+  2. 2.
+    UniProt also offers similarity-based clustering of its entries at various sequence similarity thresholds of 100%, 90%, and 50%, known as UniRef clusters.
+For instance, protein sequences that share a minimum of 50% pairwise sequence similarity will be assigned the same UniRef50 cluster.
+Recognizing a strong bias towards more commonly studied protein families in SwissProt (Figure B.1), we limited Pika-DS to a maximum of two most informative sequences per UniRef50 cluster using a custom algorithm (Section A.3).
+This resulted in a debiased set of over 257,000 protein sequences and 1.17 million information fields.
+  3. 3.
+    Lastly, we used GPT3.5 API to process the information fields for each protein entry using systematically optimized prompts (Sections A.4 and A.5) to create the Pika-DS’s three main components (example in Table B.2):
+    • Summary: A summary of each protein’s functional and biochemical properties, based solely on the provided information excluding the protein’s name.
+    • QAs: Several diverse QA pairs for each information field, formatted for LLM training.
+    • Metrics: Single-word answers to a set of predefined scientific questions serving as the ground-truth for our Biochem-Lite benchmarks (Section 3.2.2).
+A.2 Data collection from UniProt
 We retrieved XML files of all SwissProt entries longer than 30 amino acids, with the cut-off date of 14-08-2023, using UniProt’s API.
@@
 This selection, guided by an expert biochemist, included sequence (molecular mass and length), organism (top three taxonomic levels), catalytic activity (including EC number), biophysicochemical properties (pH and temperature dependence), cofactor, subunit (excluding fields containing “interact”, “associate”, or “complex”), subcellular location (excluding isoforms), and functional domains: GO (only molecular function, omitting biological process and cellular component), Gene3D, and SUPFAM.
-A.2 Debiasing SwissProt Entries Using UniRef50 Clusters
+A.3 Debiasing SwissProt Entries Using UniRef50 Clusters
 To mitigate bias in SwissProt entries, we employed UniRef50 clusters, sourced from the UniProt FTP’s idmapping file.
@@
 This debiasing process was designed to refine the Pika-DS by leveraging UniRef50 clusters, enhancing the dataset’s quality and representativeness for downstream applications.
-A.3 Example Instructions to GPT3.5
-You will receive details about a specific protein.
-Perform the following tasks: 1) Provide a factual summary, with a maximum of 500 words, that accurately and scientifically describes the functional, biochemical and structural properties of this protein based only on the provided information.
+A.4 GPT3.5 Prompt Optimization Strategy
+To ensure the high quality of GPT3.5 generated information and QA pairs, we performed a systematic prompt optimization focusing on GPT3.5’s adherence to using only the given data, summarizing complex biological reactions accurately, and avoiding speculative or ambiguous language.
+This was performed in an iterative process where at each iteration, GPT3.5 was provided with the prompt and all the extracted information for 50 randomly selected protein sequences.
+Next, 100 outputs from each of the summary statements, QA pairs and evaluation metrics were evaluated against the input information of the respective proteins by an expert biochemist and the number of incorrect, poor quality and irrelevant (correct but unrelated to the protein’s function) statements or QA pairs were noted.
+At each step, the prompt was modified to address the most problematic issues, with a target of no more than one incorrect, two poor quality and two irrelevant values in each category (optimised prompt in Section A.5).
+A.5 GPT3.5 Prompts Used for Creation of Pika-DS
+Summarising and QAs:
+The following instructions were used to create summary statements and QA annotations based on information fields collected from SwissProt:
+You will receive details about a specific protein. Perform the following tasks and print each result in a new line:
+1) Provide a factual summary, without using the protein’s name with a maximum of 500 words.
+Your summary must accurately and scientifically describe the functional, biochemical and structural properties of this protein based only on the provided information.
 Ensure that the summary follows a natural and scientific flow, starting with general information such as structure, localization and taxonomy before detailing functional and biochemical properties.
-Ensure that all key points are covered and DON’T provide any further information than what is stated in the input.
+Ensure that all key points are covered and DON’T provide any extra information than what is stated in the input.
 2) For each type of information provided, create a question-and-answer pair to elucidate an aspect of the protein’s functionality or biochemical properties without using the protein’s name.
-Use the below ; separated two-element tuples structure for the output: (summary: this protein…); (what is question 1?, the answer 1); (what is question 2?, the answer 2); …
-For both tasks if the input contains large group of properties only provide the canonical and crucial information rather than enumerating every single entry.
+Phrase your questions and answers such that they will be suitable for training a language model.
+DON’T enumerate or label the questions and print each question and its answer pair in the same line.
+- For all tasks if the input contains large group of properties only provide the canonical and crucial information rather than enumerating every single entry.
+- Where applicable, summarise enzymatic reactions into one or two of the generic classes of the activities.
 - DON’T use any of your knowledge to add additional context or information.
@@
 - AVOID using generic phrases or embellished terms such as ’highly’, ’several’, ’diverse’ and ’various’.
-A.4 Biochem-Lite Metric Questions and Motivation
+- Exactly follow the output format provided below to ensure consistency.
+Final output format: summary: [YOUR SUMMARY] QA pairs: 1) [YOUR QUESTION] [YOUR ANSWER] 2) [YOUR QUESTION] [YOUR ANSWER] […] n) [YOUR QUESTION] [YOUR ANSWER]
+Biochem-Lite Ground truth:
+The following instructions were used to collect ground-truth values for Biochem-Lite questions based on summary statements and information fields:
+You will receive details about a specific protein.
+Provide a single word answer to the following questions.
+Print each question and your answer in the same new line.
+If the question does not apply to the protein, ignore the question.
+1) Is this protein localized to the cell membrane?
+2) Is this a membrane protein?
+3) Is this protein localized to nucleus?
+4) Is this protein localized to mitochondria?
+5) Does this protein bind to DNA?
+6) Does this protein bind to RNA?
+7) Is this protein an enzyme?
+8) What are co-factors of this protein as a comma separated list?
+- Exactly follow the output format provided below to ensure consistency.
+Final output format: 1) [MY QUESTION] [SINGLE WORD ANSWER e.g.
+YES/NO/UNKNOWN] 2) [MY QUESTION] [SINGLE WORD ANSWER e.g.
+YES/NO/UNKNOWN] […] 7) [MY QUESTION] [SINGLE WORD ANSWER e.g.
+YES/NO/UNKNOWN] 8) [MY QUESTION] [Comma separated list of co-factors e.g.
+FAD,FMN/UNKNOWN]
+A.6 Biochem-Lite Metric Questions and Motivation
 mw MALE
@@
 While the information extraction process is similar, the LLM needs to perform reasoning to find the answer.
-A.5 MLP Baseline
-We use a 3-layer MLP with GELU activation on protein sequence embeddings from the Pre-trained PLM.
-Where possible, we convert each light-weight benchmarking question into a classification or value prediction task (Table 2).
-The MLP is then trained to predict the correct class labels or values using cross entropy or mean squared error loss, respectively.
-This baseline sets an upper limit for the information content inherent in PLM embeddings, with any improvements potentially indicating generalization due to the knowledge encapsulated in the LLM.
-A.6 Hyperparameter Optimization
+A.7 Metric equations used in Biochem-Lite
+  • mw MALE: The mean-absolute log error (MALE) of the predicted molecular weight (MW) is:
+    $MALE=\frac{1}{N}\sum_{i=1}^{N}|\log_{10}(\hat{MW}_{i}/MW_{i})|$ (1)
+    where $\hat{MW}_{i}$ is the predicted MW, $MW_{i}$ is the ground truth MW, and $N$ is number of examples.
+  • exact cofactor: The score is computed as:
+    $\text{Score}_{\text{exact}}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{1}(\exists w\in R_{i}:w\in GT_{i})$ (2)
+    where $R_{i}$ is the set of words in response $i$, $GT_{i}$ is the set of ground truth cofactor words for protein $i$, and $\mathbb{1}$ is the indicator function that is 1 for exactly one match and 0 otherwise.
+  • location F1: The F1 score is the harmonic mean of Precision and Recall.
+The predicted class assignments for the calculation of Precision and Recall are based on the exclusive presence of correct labels (membrane, nucleus, or mitochondrion) in the generated response.
+Responses lacking or containing multiple labels receive the none class label.
+  • is_enzyme F1: The F1 score is computed similar to location F1.
+Responses are classified as True or False based on the exclusive presence of yes or no, respectively.
+Any deviation is labeled as none.
+  • binary Q mean F1: We have:
+    $F1_{\text{binary Q mean}}=\frac{1}{3}\sum_{q=1}^{3}F1_{q}$ (3)
+    where $F1_{q}$ is the F1 score for each binary question (class assignments similar to is_enzyme F1).
+A.8 MLP Baseline
+We use a 3-layer MLP with GELU activation on protein sequence embeddings from the Pre-trained ESM2.
+Where possible, we convert each light-weight benchmarking question into a classification or value prediction task.
+More specifically we performed regression for log values of mw for mw MALE and performed classification for is_enzyme and location questions.
+The MLP is then trained to predict the value or correct class labels using mean squared error or cross entropy loss, respectively.
+This baseline sets an upper limit for the information content inherent in ESM2 embeddings.
+A.9 Training
+We devised a simple training strategy for PQA models, keeping both the PLM and LLM frozen and optimizing for the causal language model loss with AdamW.
+Hyperparameters were optimised following a greedy search as detailed in section A.10.
+All training were performed on a single A100-80GB or H100-80GB Nvidia GPU.
+Unless otherwise stated, both the QA and summary statement section of the Pika-DS was used for the training of all Pika models.
+Examples were split to train, validation and test set based on the UniRef50 cluster or EvoGroup of their respective protein sequences in 94.5%, 0.05% and 5% ratios, respectively.
+A.10 Hyperparameter Optimization
 Greedy hyper parameter search was performed for both architectures, monitoring the is_enzyme metric.
@@
 The final set of hyper parameters were as follows: learning rate: 1e-4, weight decay: 1e-4, batch size: 8, Perceiver latent size: 100, and number of Perceiver layers: 1 (Cross-Pika architecture), 4 (Self-Pika architecture).
-Appendix B Supplementary Figures and Tables
+A.11 Creation of EvoGroups
+EvoGroups were decided by starting from a random sequence in the data, identifying all its related sequences using JackHammer (Potter et al., 2018) with a very lenient cut-off threshold for e-E-score of 1.0 to ensure a broad grouping of related proteins.
+For each randmoly selected sequence, all related sequences were marked as belonging to the same EvoGroup and were removed from the remainder of the dataset for further iterations with newly selected sequences.
+The process was repeated until all sequences were assigned an EvoGroup.
+Appendix B Supplementary Figures & Tables
+Figure B.1: Over-representation bias in SwissProt database.
+Figure B.2: Comparison of sequence bias in SwissProt database, before and after filtering.
+Members of the top100 largest UniRef50 clusters are colored.
+The strong overrepresentation of highly studied protein groups is apparent in the unfiltered plot, while the Pika-DS shows a significant reduction in this bias
+Figure B.3: Cross-correlation of Biochem-ReAct Benchmarks with Biochem-Lite/Linguistic Benchmarks. Absolute values.
+# total # unique # words # tokens
+Proteins 257 K 252 K - 105.3 M
+Info fields 1.17 M 319 K 10.8 M 29.5 M
+Summaries 1.33 M 861 K 22.5 M 36.4 M
+Questions 1.71 M 98 K 15.7 M 14.9 M
+Answers 1.71 M 783 K 18.4 M 31.8 M
+Table B.1: Pika-DS size and token counts. Values for Summaries represent the number of sentences.
 uniref UniRef50_Q9HUN4
@@
 Table B.2: An example entry in Pika-DS representing Uniprto ID A4VQM5
-Figure B.1: Over-representation bias in SwissProt database.
-Figure B.2: Comparison of sequence bias in SwissProt database, before and after filtering.
-Members of the top100 largest UniRef50 clusters are colored.
-The strong overrepresentation of highly studied protein groups is apparent in the unfiltered plot, while the Pika-DS shows a significant reduction in this bias
-Figure B.3: Evaluating the Effectiveness of Biochem-Lite Versus Traditional Linguistic Metrics for Scientific Accuracy of PQA.
-Statistical significance is determined through a one-tailed paired t-test across three randomly seeded data subsets and model training (significance guide: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$).
-By design, Self-Pika is expected to outperform Pika w/o PLM across all metrics, except for the mw MALE metric.
-Figure B.4: Cross-correlation of Biochem-ReAct Benchmarks with Biochem-Lite/Linguistic Benchmarks. Absolute values.
+Random
+LLM
+only
+self-Pika
+w/o PLM
+self-Pika
+F1 Score 0.49 0.00 0.43 0.99
+Accuracy 0.50 0.00 0.51 0.99
+Table B.3: Performance on Control Question.